### v1

# RuleR: Improving LLM Controllability by Rule-based Data Recycling

Abstract
Large language models (LLMs) still lack delicate controllability over their responses, which is critical to enhancing their performance and the user experience. However, curating supervised fine-tuning (SFT) datasets to improve LLM controllability usually relies on human experts or proprietary LLMs, which requires additional costs. To bridge this gap, we propose Rule-based Data Recycling (RuleR), a data augmentation method incorporating multiple constraints into the original data samples according to predefined rules, which creates new training tasks to consolidate the controllability of LLMs. Instead of creating new data from scratch, RuleR “recycles” existing data by simply applying rule-based edits to their responses and appending the rule-instructions in their original instructions. Experimental results demonstrate RuleR’s effectiveness in improving LLM controllability while maintaining general instruction-following capabilities. The code will be released on https://github.com/MingLiiii/RuleR.
RuleR: Improving LLM Controllability by Rule-based Data Recycling
Ming Li*1, Han Chenthanks: Equal Contribution., Chenguang Wang*2, Dang Nguyen1, Dianqi Li, Tianyi Zhou1 1University of Maryland 2Stony Brook University {minglii, tianyi}@umd.edu
1 Introduction
Figure 1: Comparing existing methods (top) and our RuleR (bottom) for enhancing LLM controllability. Most existing methods rely on extra-human/model supervision to generate or edit instructions and responses, neglecting the remaining potential of the original data. On the contrary, RuleR demonstrates that simple rule-based (human/model-free) editing of existing data can greatly improve LLM controllability.
Despite the remarkable advancement of the current LLMs and the continuous efforts to build high-quality supervised fine-tuning (SFT) datasets Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a, b); Jiang et al. (2023), one critical aspect is how these LLMs better interact with human responses to maximize their utility and effectiveness for end-users Liu et al. (2024). According to the systematic investigation from Liu et al. (2024), it is essential for LLMs to constrain their outputs to follow specific user-defined formats or standards. Without constraints on responses, it is hard to directly put the free-formed responses into practical usage. This is also verified in the area of LLM Agents Li et al. (2024f); Chen et al. (2023, 2024b); Zhang et al. (2024), in which a predefined format is always necessary.
However, existing methods for building SFT datasets focus mainly on general instructions Wei et al. (2022); Wang et al. (2022); Taori et al. (2023); Xu et al. (2023); Zhou et al. (2023a); Li et al. (2023a); Zhang et al. (2023); Xu et al. (2024), neglecting the potential user-specific constraints and thus lacking delicate controllability of the response on user-desired lengths or formats Chen et al. (2024a); Xia et al. (2024). One potential method to enhance existing SFT data with multi-constraint controllability is to rewrite or modify instructions and responses by expert humans/models Xu et al. (2023); Li et al. (2023a, 2024b, 2024a); He et al. (2024) as shown in Figure 1 (top). However, these methods directly discard or rewrite the original data. Their data curation is costly and inefficient, requiring careful editing by human experts or proprietary LLMs. This raises the question: Can we “recycle” existing SFT datasets for controllability tuning by enforcing various types of constraints to them without human/model editing?
Drawing inspiration from recent work like Mosaic-IT Li et al. (2024c), which develops a human/model-free method augmenting SFT data, and IFEval Zhou et al. (2023b), which utilizes verifiable constraints to evaluate LLMs’ controllability, we propose Rule-based Data Recycling (RuleR), which automatically edits existing SFT data for improving LLM controllability. The key insight of RuleR is to replace human/model editing with various predefined rules of constraining patterns. Specifically, we manually inspect and construct several verifiable constraint rules, in which the original instructions are appended with various constraints, and the responses are optionally modified to fulfill these constraints based on their original characteristics. Our predefined rules cover a wide range of diverse constraints ranging from high-level constraints, e.g., controlling the word frequency in the response, to lower-level constraints, e.g., setting specific wrapping formats [^1] The complete list of rules and descriptions can be found in the Appendix C. . Each rule contains (1) a set of constraint templates that serve as additional rule-instructions, on constraints the LLM should follow, and (2) specified operations that alternately edit the instruction and response to reach an alignment between them (i.e., all the constraints in the instruction hold in the response). For each sample from the original dataset, we randomly draw several rules to be applied in the editing. This produces an augmented sample with the constraints enforced that can be used for controllability tuning.
An example is shown in Figure 1 (bottom). The original instruction is “What type of fruit would be a good addition to a fruit salad?”, and the corresponding response is “Grapes, kiwis, oranges, apples.”. If the constraint “Num of Nouns” is sampled and implemented on this sample, the augmented instruction becomes “What type of fruit would be a good addition to a fruit salad? Ensure the response has more than 3 nouns.”, and the original response remains unchanged on this type of rule. In this example, the constraint (number of nouns) is directly decided by the original data sample (4 nouns), thus avoiding the regeneration of response.
Extensive experimental results on the IFEval (Instruction-Following Eval) Zhou et al. (2023b) benchmark on various datasets demonstrate the effectiveness of our human/model-free method in enhancing LLM controllability. Moreover, the instruction-following metrics also show that RuleR still preserves the general instruction-following ability promoted by the original SFT dataset. To the best of our knowledge, RuleR is the first human/model-free data augmentation approach designed to improve LLM controllability in enforcing multiple constraints to LLM-generated response.
2 Method
2.1 Preliminaries
Given a supervised finetuning dataset $D$, there are $n$ data samples each of which is represented by a tuple $(x_{i},y_{i})$, where $x_{i}$ represents the instruction and $y_{i}$ represents the corresponding response. Let $p_{\theta}(\cdot)$ denote the LLMs to be trained, with parameters $\theta$. In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by maximizing the following objective on each data $(x_{i},y_{i})$, $y_{i,j}$ represents the $j_{th}$ token of response $y_{i}$ and $l_{i}$ represents the token length:
$\max_{\theta}\sum_{j=1}^{l_{i}}\log p_{\theta}\left(y_{i,j}|x_{i}\right),$ (1)
2.2 Rule-based Data Recycling
The basis for augmenting existing SFT datasets is to reconstruct instructions and responses. Most existing methods utilize human experts or strong teacher models for this process, which is costly, inefficient, and fails to get fully used to the existing data, as represented in Figure 1 (top). Inspired by Mosaic-IT which proposes a human/model-free augmentation method to reconstruct instruction-response pairs, we are motivated to design a rule-based method for controllability augmentation.
Model Dataset Method Instruction Following Eval Pair-wise Open LLM Leaderboard
Prompt (S) Inst (S) Prompt (L) Inst (L) Winning Score Average A H M T
Mistral-7B Alpaca-GPT4 Baseline 32.72 42.45 35.67 45.44 1.000 61.24 56.23 81.07 56.22 51.42
Ours 39.56 51.44 43.44 55.40 1.044 61.05 56.66 80.50 57.24 49.81
Alpaca Baseline 33.64 44.60 36.23 47.96 1.000 55.15 51.96 74.61 52.85 41.20
Ours 35.12 46.76 37.89 49.52 1.158 56.21 54.61 77.70 54.54 38.00
Wizard-70k Baseline 37.34 48.68 40.85 52.04 1.000 59.38 54.61 79.96 55.68 47.27
Ours 46.77 57.07 49.17 59.71 1.168 59.75 55.38 80.75 55.59 47.27
Recycled Wizard Baseline 30.87 42.41 35.86 46.40 1.000 59.61 54.10 77.85 57.61 48.87
Ours 39.56 51.08 45.10 56.24 0.987 60.43 56.66 78.01 58.61 48.43
DEITA 6K Baseline 41.22 51.08 44.55 54.92 1.000 64.82 60.41 82.52 61.57 54.76
Ours 45.10 54.32 47.69 56.95 1.010 65.34 61.16 82.78 62.74 54.66
Llama2-7B Alpaca-GPT4 Baseline 26.25 36.33 30.31 40.29 1.000 58.71 54.69 80.05 47.89 52.21
Ours 32.35 42.09 35.30 45.56 1.070 59.77 56.74 80.67 48.45 53.21
Alpaca Baseline 31.42 40.77 33.46 43.17 1.000 55.25 54.35 78.65 47.02 40.98
Ours 34.38 44.36 37.34 47.36 1.023 55.24 54.61 78.76 46.17 41.42
Wizard-70k Baseline 31.24 44.24 35.49 48.68 1.000 57.09 54.18 79.25 46.93 48.02
Ours 38.82 50.12 42.33 53.48 1.087 57.25 55.20 79.81 46.61 47.38
Llama2-13B Alpaca-GPT4 Baseline 32.90 44.60 36.23 48.08 1.000 61.47 58.70 83.12 54.13 49.92
Ours 36.60 47.00 37.89 49.28 0.977 61.96 59.47 82.88 53.98 51.52
Alpaca Baseline 34.94 44.36 36.41 46.29 1.000 57.63 57.25 81.23 54.13 37.91
Ours 36.04 48.20 41.22 52.88 0.977 57.16 57.17 81.11 52.70 37.65
Wizard-70k Baseline 43.07 53.84 46.40 57.67 1.000 61.24 57.04 83.39 55.76 48.78
Ours 45.47 58.15 50.09 61.99 1.010 60.84 58.28 82.37 54.35 48.36
Table 1: The comparisons on the Instruction Following Eval, Pair-wise Comparison Winning Score, and the Open LLM Leaderboard, on 3 different base models and several different instruction tuning datasets. The baselines are models trained with the original dataset. Prompt and Inst represent Prompt-level and Instruction-level accuracy; S and L represent Strict and Loose versions. A, H, M, and T represent ARC, HellaSwag, MMLU, and TruthfulQA.
With this goal, the first challenge is the reformulation of instructions. Due to the lack of human experts or LLMs to generate the corresponding responses for any given instructions, it is not feasible to add arbitrary constraints to the original instructions. To solve this problem, we propose to only incorporate constraints that are compatible with the original data sample. Specifically, we propose to enhance LLM controllability by Rule-based Data Recycling (RuleR), in which the rules and constraints used to reformulate instruction are determined by the original data, ensuring the alignment between instruction and response. Motivated by IFEval, we manually inspect and propose several verifiable rules in which the responses are not required for regeneration by LLMs or human experts, but only need optional slight adjustments. The complete list of rules is listed in Table 3 ranging from high-level constraints like term frequency to lower-level constraints like specific wrapping formats.
During implementation, we instantiate each predefined rule as a tuple , $(\mathbf{S}_{k},f_{k},g_{k})$, where $\mathbf{S}_{k}$ represents the set of manually curated rule-instruction templates for the $k^{th}$ rule, $f_{k}$ and $g_{k}$ represent the function to reformulate instructions and optionally modify responses respectively, as shown in Figure 1(bottom). For each data sample $(x_{i},y_{i})$, the augmented instruction $x_{i,aug}$ will be reformulated according to the characteristics of the original data and the corresponding template sets:
$x_{i,aug}=f_{k}(x_{i},y_{i},\mathbf{S}_{k})$
where specifically, a random rule-instruction template will be sampled and filled with the corresponding characteristics as the additional constraint to the original instruction. Then the rule-instruction is concatenated with the original instruction becoming the augmented instruction. Similarly, the augmented response $y_{i,aug}$ will be optionally modified based on the selected rule $k$:
$y_{i,aug}=g_{k}(x_{i},y_{i},\mathbf{S}_{k}),$
where the function $g_{k}$ varies dramatically for different rules: For some of the rules, no modification needs to be done to the original response, while for others, some arbitrary part of the response needs to be replaced with the desired user format.
3 Experiments
3.1 Experimental Setup
We conduct experiments on several datasets including the Alpaca Taori et al. (2023), Alpaca-GPT4 Peng et al. (2023), WizardLM Xu et al. (2023), Recycled WizardLM Li et al. (2023a), DEITA Liu et al. (2023), and on several base LLMs including Llama2-7B, Llama2-13B Touvron et al. (2023b), and Mistral-7B Jiang et al. (2023), introduction on datasets can be found in the Appendix A. We utilize the prompt and code base from Vicuna Chiang et al. (2023a) and flash attention Dao et al. (2022). The Adam optimizer Kingma and Ba (2017) is utilized with the batch size of $128$ and with the max token length of $2048$. For training on Llama2-7B and Llama2-13B Touvron et al. (2023b), the maximum learning rate is set to $2\times 10^{-5}$ with a warmup rate of $0.03$ for $3$ epochs. For training on Mistral-7B Jiang et al. (2023), the maximum learning rate is set to $1\times 10^{-5}$ with a warmup rate of $0.1$ for $2$ epochs. When utilizing our method, we run the augmentation process $3$/$2$ times to simulate the epochs of training. These augmented data are then mixed together and used for training $1$ epoch. All other configurations are kept the same as the baselines.
3.2 Experimental Results
The main experimental results are presented in Table 1, containing the performance comparison on the Instruction Following Eval Zhou et al. (2023b), Pair-wise Comparison Winning Score, and the Open LLM Leaderboard Gao et al. (2021a), on 3 different base models and several different instruction tuning datasets. The Pair-wise Comparison Winning Score is calculated as (Num(Win)--Num(Lose))///Num(All) $+1$ and the values that are greater than $1.0$ represent better responses generated. Detailed descriptions of evaluation metrics can be found in the Appendix B.
Compared to the baseline, our method has consistent improvements on the IF Eval benchmark, across different base models and datasets, which aims at measuring LLMs’ constraint-following abilities by using verifiable instructions. It is astonishing that our method can improve the IF Eval scores by approximately $10\%$ on some of the configurations, by just utilizing the rule-based recycling method on the original data, without any human/model edition. Moreover, the performances keep being positive on originally diverse and high-quality datasets like Recycled WizardLM Li et al. (2023a) and DEITA Liu et al. (2023), which further verify the potential of our method. Compared with existing methods which enhance LLMs’ constraint controllability by generating totally new data, our method focuses more on fully utilizing the potential of existing data.
Furthermore, our method not only improves the constraint controllability but also keeps the general instruction-following ability of the original data. The Pair-wise comparison and Open LLM leaderboard results showcase comparable or sometimes better performances compared with the baseline models. We hypothesize that the additional constraints might complicate the original instructions, thus leading to potentially improved instruction-following abilities.
3.3 Ablation Studies
Evaluation Metrics IF Eval Pair-wise Open LLM
Baseline 39.07 1.000 61.24
Single Temp. 42.62 0.987 61.43
Diverse Temp. (*) 47.46 1.044 61.05
Max Rule = 1 46.14 1.168 61.22
Max Rule = 2 47.09 1.117 61.15
Max Rule = 3 (*) 47.46 1.044 61.05
Max Rule = 4 46.89 1.003 61.55
Max Rule = 5 44.36 1.013 60.06
Aug Rate = 0.1 41.72 1.020 61.34
Aug Rate = 0.3 42.08 1.037 61.20
Aug Rate = 0.5 46.55 1.111 61.31
Aug Rate = 0.7 46.23 1.111 61.18
Aug Rate = 0.9 (*) 47.46 1.044 61.05
Table 2: Ablations. The experiments are conducted on Mistral-7B on Alpaca-GPT4 data and the configuration notated with “(*)” represents our default setting. “IF Eval” and “Open LLM” represent the averaged scores.
In this section, ablation experiments are conducted on Mistral-7B with the Alpaca-GPT4 dataset.
As shown in Table 2, “Single Temp.” represents utilizing only one rule-insturtcion template for each rule, while “Diverse Temp.” represents utilizing several different templates (with the same meaning) for each rule and those templates are randomly sampled during the augmentation. The superiority of “Diverse Temp.” showcases the necessity of providing diverse templates to avoid potential overfitting. “Max Rule = x” represents the setting in which at most x different rules can be sampled and utilized on each original data sample. The performances of models, whose Max Rule is set under 5, are promising, while the performance declines on all 3 evaluation metrics when the Max Rule is set as 5. These results indicate that too many rules utilized together in one sample might be harmful to LLM capability. “Aug Rate = x” represents there is a probability of x to apply our augmentation to each sample. It can be observed that increasing the augmentation rate is beneficial to the controllability of LLMs, while slightly harmful to the general instruction-following capability. However, the gaps on IF Eval are much larger than the other 2 metrics, indicating this rate will mostly influence the multi-constraint controllability of LLM.
4 Conclusion
In this work, we proposed a novel method, Rule-based Data Recycling (RuleR) which utilizes the constraints derived from the existing data to enhance the dataset and incorporates these constraints into the training process, thereby improving the multi-constraint controllability of LLMs. This approach presents a promising direction for developing SFT-augmenting methods to improve LLM controllability, paving the way for more efficient advancements of LLM.
Limitations
Our method focuses on improving LLM controllability by rule-based data augmentation to avoid extra regeneration processes from humans or models. Though the alleviation of reliance on human/model editing is our main contribution, it inevitably limits the types of constraints that can be implemented on the original data sample. Currently, all the constraint rules are based on verifiable shallow syntactic characteristics like the occurrence and frequency of words or sentences, etc, lacking constraint rules that focus more on the semantic feature or content. Thus, our method shows great potential to be further enhanced if further involves small models, which will still remain efficient and avoid the regeneration process.
Appendix A Training Dataset
We utilize 5 SFT datasets to evaluate the effectiveness of our method:
Alpaca dataset Taori et al. (2023): This dataset consists of 52,000 instruction-following samples created using the self-instruct paradigm Wang et al. (2023b) and OpenAI’s text-davinci-003 model. Characterized as a classical dataset with moderate-quality attributes, it serves as the fundamental validation.
Alpaca-GPT4 dataset Peng et al. (2023): This dataset is an enhanced Alpaca dataset that includes responses generated by GPT-4.
WizardLM dataset Xu et al. (2023): This dataset is generated by the novel Evol-Instruct method, which utilizes ChatGPT-3.5 to rewrite instructions step by step into more complex ones and generate the corresponding responses. We utilize the 70k version in our method, which comprises 70,000 high-quality SFT samples.
Recycled WizardLM Dataset Li et al. (2023a): This dataset is an improved version of the WizardLM dataset, by utilizing the Reflection-Tuning method. In the Reflection-Tuning, the initial dataset undergoes two main phases: Reflection on Instruction and Reflection on Response. In the first phase, specific criteria are carefully curated to evaluate and refine the initial instructions. During the second phase, responses are thoroughly examined and improved to align with the refined instructions. This process generates a dataset with superior quality compared to the original dataset.
DEITA dataset Liu et al. (2023): This dataset leverages the DEITA (Data-Efficient Instruction Tuning for Alignment) method to select high-quality data from a pool comprised of several high-quality datasets, such as WizardLM and Alpaca. DEITA employs a score-first, diversity-aware data selection strategy to optimize the selection process. This strategy uses a GPT-as-a-judge scoring system that combines complexity and quality in a practical and straightforward manner. The scores are incorporated with the diversity-based selection, ensuring that all the data maintains high standards of complexity, quality, and diversity.
Appendix B Evaluation Metrics
We employ 3 commonly accepted metrics for the evaluation, including IFEval (Instruction-Following Eval), Pair-wise Comparison, and Open LLM Leaderboard
IFEval Zhou et al. (2023b) is the primary evaluation metric employed in our study due to its compatibility with our motivation. It focuses on evaluating how LLMs follow various additional constraints, such as specifying a word count or requiring the inclusion of certain keywords a specified number of times. To avoid the utilization of LLMs during evaluation, it proposes 25 distinct types of verifiable instructions. There are 541 prompts in total and each of them incorporates one or more of these verifiable instructions, ensuring the comprehensiveness of the evaluation. IFEval serves as a great inspiration for our method, and there exist semantical overlappings between their verifiable instructions and our rules. However, the specific prompts used in IFEval are kept unknown in the construction of our rule templates, avoiding potential data leakage. Moreover, IFEval only needs to verify the responses, while our method needs to modify the responses for the training, which pushes this process a step further. Consequently, this benchmark not only facilitates a comprehensive comparison but also provides valuable insights that align with our purpose.
Pair-wise Comparison involves evaluating responses from LLMs like GPT-4, especially in open-domain contexts. This method has shown a notable alignment with human assessments, providing a credible evaluative foundation Zheng et al. (2024); Li et al. (2023b); Sottana et al. (2023). We utilize test instruction sets from WizardLM Xu et al. (2023) and Vicuna Chiang et al. (2023b), comprising 218 and 80 diverse human-curated instructions, respectively. Following the framework by Li et al. (2024e, d), we prompt judging LLM to rate responses on a scale from 1 to 10 across multiple dimensions such as relevance and accuracy. To mitigate positional bias Ko et al. (2020); Wang et al. (2023a), responses are presented in two distinct sequences. Evaluation outcomes are categorized into the ’win-tie-lose’ system for each instruction. One model’s response can be noted as Win/Lose only if it has a Higher/Lower score in both directions. The specific evaluation prompt is presented in Figure 2.
Open LLM Leaderboard employs Eval Harness Gao et al. (2021a) to systematically assess generative language models through diverse tasks. Key benchmarks include ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2020), and TruthfulQA Lin et al. (2021), collectively evaluating reasoning abilities, common-sense knowledge, and factual accuracy. This approach highlights LLM strengths and areas for improvement in critical aspects of language understanding and generation Gao et al. (2021b).
  Prompt for Performance Evaluation
  System Prompt
You are a helpful and precise assistant for checking the quality of the answer.
User Prompt
[Question]
Question
[The Start of Assistant 2’s Answer]
Answer 2
[The End of Assistant 2’s Answer]
[The Start of Assistant 2’s Answer]
Answer 2
[The End of Assistant 2’s Answer]
We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above.
Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance.
Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.
Figure 2: The prompt we used to request GPT4 to evaluate the responses.
Appendix C Predefined Rules
Rule Type Rule Name Example Rule-Instruction Templates
Keyword Frequency Keyword Appearance Ensure {Keyword} is in the response.
Keyword Frequency Keyword Frequency Ensure there are more/less/exact {N} {Keyword} in the response.
Number Constraint Num of Adjectives Ensure the response has more/less/exact {N} adjectives.
Number Constraint Num of Nouns Ensure the response has more/less/exact {N} nouns.
Number Constraint Num of Verbs Ensure the response has more/less/exact {N} verbs.
Number Constraint Num of Characters Ensure the response has more/less/exact {N} characters.
Number Constraint Num of Letters Ensure the response has more/less/exact {N} letters.
Number Constraint Num of Words Ensure the response has more/less/exact {N} words.
Number Constraint Num of Sentences Ensure the response has more/less/exact {N} sentences.
Number Constraint Num of Paragraphs Ensure the response has more/less/exact {N} paragraphs.
Number Constraint Num of Bullets Ensure the response has more/less/exact {N} bullet points.
Repetition Instruction Repetition Repeat the instruction before providing the response.
Repetition Response Repetition Repeat the response {N} times.
Case All Up Case Ensure the response is all in upper case.
Case All Low Case Ensure the response is all in lowercase.
Case Target Letter Case Ensure all the letters {x} in the response are in uppercase.
Case Target Keyword Case Ensure all the word {Keyword} in the response are in uppercase.
Case Target Sentence Case Ensure {i}-th sentence in the response is in uppercase.
Case Target Paragraph Case Ensure {i}-th paragraph in the response is in uppercase.
Punctuation All All Removal Ignore all punctuation in the response.
Punctuation All All Replacement Use {Symbol} to replace all punctuation in the response.
Punctuation Target Target Removal Ignore {Punctuation} punctuation in the response.
Punctuation Target Target Replacement Use {Symbol} to replace {Punctuation} in the response.
Format Wrapping Keyword Wrapping Ensure every {Keyword} is wrapped in {Format} in the response.
Format Wrapping Sentence Wrapping Ensure {i}-th sentence is wrapped in {Format} in the response.
Format Wrapping Bullet Wrapping Ensure {i}-th bullet point is wrapped in {Format} in the response.
Format Wrapping Paragraph Wrapping Ensure {i}-th paragraph is wrapped in {Format} in the response.
Formatted Repeating Instruction Wrapping Repeat the instruction in {Format} before providing the response.
Formatted Repeating Response Wrapping Repeat the response {N} times in {Format}.
Table 3: The list of predefined constraint rules.
In this section, we will dive into the predefined rules describing each constraint specifically.
Keyword Appearance simulates the scenario where specific keywords are required to appear in the responses. In this rule, several non-stop words are randomly selected from the original data sample and used as the desired characteristics. The placeholder Keyword in the constraint template will be replaced by the sampled keyword as the rule-instruction. This process can be repeated to simulate the constraints on multiple keywords. The augmented instruction is the concatenation of the original instruction and rule-instruction. The original response does not need to be modified in this rule and is directly used as the augmented response.
Keyword Frequency simulates controlling the frequency of specific keywords in generated responses. In this rule, several non-stop words and their frequencies are randomly sampled and used as the desired characteristics. There are three random sub-situations in the rule: More, Less, or Equal. In the “Equal” situation, the placeholders {N} and {Keyword} will be directly replaced by the sampled keyword and its frequency. In the “More” or “Less” situations, a small random number $x$ will be randomly generated to adjust the keyword frequency to meet the desired constraint template, such as “Ensure there are more than {N - x} {Keyword}” or “Ensure there are fewer than {N + x} {Keyword}.” This process can be repeated to simulate constraints on multiple keywords. The augmented instruction is the concatenation of the original instruction and rule-instruction. The original response does not need to be modified in this rule and is directly used as the augmented response.
Num of Adjectives simulates controlling the total number of adjectives in generated responses. In this rule, the adjectives in the original response are identified and counted using part-of-speech tagging (POS). There are three possible sub-situations in the rule: More, Less, or Exact. In the “Exact” situation, the placeholder {N} will be replaced by the number of adjectives. In the “More” or “Less” situations, a small random number $x$ is randomly generated to adjust N to meet the constraints, such as “Ensure the response has more than {N - x} adjectives” or “Ensure the response has fewer than {N + x} adjectives.” This process can only be used once for each sample. The augmented instruction is the concatenation of the original instruction and rule-instruction. The original response does not need to be modified in this rule and is directly used as the augmented response.
Num of Nouns simulates controlling the number of nouns in generated responses, similar to the “Num of Adjectives”.
Num of Verbs simulates controlling the number of verbs in generated responses.
Num of Characters simulates controlling the number of characters in generated responses.
Num of Letters simulates controlling the number of letters in generated responses.
Num of Words This rule simulates controlling the number of words in generated responses.
Num of Sentences simulates controlling the number of sentences in generated responses. The sentences from the original response are segmented by utilizing dependency parsing.
Num of Paragraphs simulates controlling the number of paragraphs in generated responses. The paragraphs from the original response are segmented by regular expressions.
Num of Bullets simulates controlling the number of bullet points in generated responses. The bullet points from the original response are segmented by regular expressions.
Instruction Repetition simulates the scenario where the LLM is requested to repeat the instructions before providing the response. This process can be applied only once for each instruction. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the concatenation of repeated original instruction and the response.
Response Repetition simulates the scenario where the LLM is requested to repeat the responses several times. In this rule, the response is repeated {N} times, where {N} is a random number. This process can only be applied once per data sample. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the concatenation of N identical responses.
UP Case simulates requesting the entire response is required in uppercase. In this rule, the original response is converted to uppercase format entirely. This process can only be used once for each response. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the all-uppercase version of the original response.
Low Case simulates requesting the entire response is required in lowercase.
Letter Case simulates the scenario where specific types of letters in the response are required to be in uppercase. In this rule, the specific letter x is sampled from the response, and all occurrences of this letter in the response are capitalized. This process can be repeated on different random letters. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the original response with all specific letters in uppercase.
Keyword Case simulates the scenario where specific keywords in the response are required to be in uppercase. The specific keyword Keyword is sampled from the response.
Sentence Case simulates the scenario where the specific sentences in the response are required to be in uppercase. The index of the sentence i is randomly selected within the total number of sentences in the response.
Paragraph Case simulates the scenario where the specific paragraphs in the response are required to be in uppercase. The index of the paragraph i is randomly selected within the total number of paragraphs in the response
All Removal simulates controlling LLM to ignore the use of punctuation. In this rule, the punctuation marks in the response will be removed completely, and the new response will serve as the augmented response. This process can only be used once for each sample.
All Replacement simulates controlling LLM to replace all the original punctuation with a predefined symbol {Symbol}. In this rule, the punctuation marks in the response will be replaced, and the new response will serve as the augmented response. This process can only be used once for each sample.
Target Removal simulates the scenario where a specific type of punctuation mark {Punctuation} in the response is ignored. In this rule, a random type of punctuation is identified, and all occurrences of this mark are removed, the new response will serve as the augmented response. This process can only be used once for each sample to avoid confusion.
Target Replacement simulates the scenario where a specific type of punctuation mark {Punctuation} in the response is replaced by a specified symbol {Symbol}. In this rule, a random type of punctuation mark is identified, and all occurrences of this mark will be replaced by a predefined symbol in the response. This process can only be used once.
Keyword Wrapping simulates the scenario where specific keywords in the response are required to be wrapped in a specified format. In this rule, a randomly chosen {Keyword} is identified, and all occurrences of this keyword are wrapped in the randomly specified {Format} in the response. This process can be repeated several times on different words with various formats. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the original response with all keywords wrapped in the format.
Sentence Wrapping simulates the scenario where specific sentences in the response are required to be wrapped in a specified format. The index of the sentence i is randomly selected within the total number of sentences in the response.
Bullet Wrapping simulates the scenario where specific bullet points in the response are required to be wrapped in a specified format. The index of the bullet point i is randomly selected within the total number of bullet points in the response.
Paragraph Wrapping simulates the scenario where a specific paragraph in the response are required to be wrapped in a specified format. The index of the paragraph i is randomly selected within the total number of paragraphs in the response.
Instruction Wrapping simulates a scenario where the original instruction is required to be repeated in a specified format before providing the response. In this rule, the original instruction is restated with wrapping in the randomly chosen {Format} before giving the actual response. This process can be applied only once for each instruction. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the concatenation of the original instruction wrapped in the specific format and the response.
Response Wrapping simulates a scenario where the response wrapped in the specific format is required to be repeated several times. In this rule, the response is repeated {N} times wrapeed in the specified {Format}, with {N} and {Format} being randomly selected. This process can be applied only once. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the concatenation of the repeated response wrapped in the format.

### v3

@@
 Abstract
-Large language models (LLMs) still lack delicate controllability over their responses, which is critical to enhancing their performance and the user experience.
-However, curating supervised fine-tuning (SFT) datasets to improve LLM controllability usually relies on human experts or proprietary LLMs, which requires additional costs.
-To bridge this gap, we propose Rule-based Data Recycling (RuleR), a data augmentation method incorporating multiple constraints into the original data samples according to predefined rules, which creates new training tasks to consolidate the controllability of LLMs.
-Instead of creating new data from scratch, RuleR “recycles” existing data by simply applying rule-based edits to their responses and appending the rule-instructions in their original instructions.
-Experimental results demonstrate RuleR’s effectiveness in improving LLM controllability while maintaining general instruction-following capabilities.
-The code will be released on https://github.com/MingLiiii/RuleR.
+Despite the remarkable advancement of Large language models (LLMs), they still lack delicate controllability under sophisticated constraints, which is critical to enhancing their response quality and the user experience.
+While conditional supervised fine-tuning (SFT) can potentially improve LLM controllability, curating new SFT data to fulfill the constraints usually relies on human experts or proprietary LLMs, which is time-consuming and expensive.
+To bridge this gap, we propose Rule-based Data Recycling (RuleR), a human/LLM-free data augmentation method incorporating multiple constraints into the original SFT data.
+Instead of creating new responses from scratch, RuleR integrates linguistic or formatting rules into the original instructions and modifies the responses to fulfill the rule-defined constraints.
+Training on the “recycled” data consolidates LLM capability to generate constrained outputs.
+Extensive experiments demonstrate RuleR’s effectiveness in improving LLM controllability while maintaining general instruction-following performance.
+RuleR’s code is released on https://github.com/tianyi-lab/RuleR.
 RuleR: Improving LLM Controllability by Rule-based Data Recycling
@@
 1 Introduction
-Figure 1: Comparing existing methods (top) and our RuleR (bottom) for enhancing LLM controllability.
-Most existing methods rely on extra-human/model supervision to generate or edit instructions and responses, neglecting the remaining potential of the original data.
-On the contrary, RuleR demonstrates that simple rule-based (human/model-free) editing of existing data can greatly improve LLM controllability.
-Despite the remarkable advancement of the current LLMs and the continuous efforts to build high-quality supervised fine-tuning (SFT) datasets Brown et al.
+Figure 1: Comparing widely-used data generation strategy (top) and RuleR (bottom) enhancing LLM controllability.
+Most existing methods rely on human/model supervision or reflection to generate new instructions and responses.
+However, discarding existing data is a waste of efforts taken to curate them by human experts or expert LLMs.
+On the contrary, our RuleR demonstrates that simple rule-based (human/model-free) editing of existing data can greatly improve LLM controllability.
+Figure 2: Examples of the data recycling workflows in Rule-based Data Recycling (RuleR).
+(a), (b) and (c) select different predefined rules to modify the original data to fulfill constraints on the complexity or format of the response.
+The differences in new responses are highlighted in red, the example in (a) has already satisfied the appended constraint, thus the response is kept unchanged.
+Despite the remarkable advancement of the current Large language models (LLMs) and the continuous efforts to build high-quality supervised fine-tuning (SFT) datasets Brown et al.
 (2020); OpenAI (2023); Touvron et al.
 (2023a, b); Jiang et al.
-(2023), one critical aspect is how these LLMs better interact with human responses to maximize their utility and effectiveness for end-users Liu et al.
+(2023), one critical challenge is to generate responses better interacting with humans, with the utility and effectiveness maximized for end-users Liu et al.
 (2024).
 According to the systematic investigation from Liu et al.
-(2024), it is essential for LLMs to constrain their outputs to follow specific user-defined formats or standards.
-Without constraints on responses, it is hard to directly put the free-formed responses into practical usage.
-This is also verified in the area of LLM Agents Li et al.
+(2024), it is essential for LLMs to constrain their outputs to follow user-specified formats or characteristics.
+In various practical applications, without any constraint or format being enforced, free-formed responses are not legal or directly applicable.
+It has also been verified on LLM Agents Li et al.
 (2024f); Chen et al.
 (2023, 2024b); Zhang et al.
-(2024), in which a predefined format is always necessary.
-However, existing methods for building SFT datasets focus mainly on general instructions Wei et al.
+(2024) that enforcing predefined formats is necessary for various tasks.
+However, existing SFT datasets are mainly composed of general instructions without user-specified constraints Wei et al.
 (2022); Wang et al.
@@
 (2023); Xu et al.
-(2024), neglecting the potential user-specific constraints and thus lacking delicate controllability of the response on user-desired lengths or formats Chen et al.
+(2024) and thus result in models lacking delicate controllability of the lengths and format of responses Chen et al.
 (2024a); Xia et al.
 (2024).
-One potential method to enhance existing SFT data with multi-constraint controllability is to rewrite or modify instructions and responses by expert humans/models Xu et al.
+To enhance the utility of existing SFT data in improving the controllability of LLMs, a potential method is to rewrite or modify instructions and responses by experts such as humans/LLMs Xu et al.
 (2023); Li et al.
 (2023a, 2024b, 2024a); He et al.
-(2024) as shown in Figure 1 (top).
-However, these methods directly discard or rewrite the original data.
-Their data curation is costly and inefficient, requiring careful editing by human experts or proprietary LLMs.
-This raises the question: Can we “recycle” existing SFT datasets for controllability tuning by enforcing various types of constraints to them without human/model editing?
-Drawing inspiration from recent work like Mosaic-IT Li et al.
-(2024c), which develops a human/model-free method augmenting SFT data, and IFEval Zhou et al.
-(2023b), which utilizes verifiable constraints to evaluate LLMs’ controllability, we propose Rule-based Data Recycling (RuleR), which automatically edits existing SFT data for improving LLM controllability.
-The key insight of RuleR is to replace human/model editing with various predefined rules of constraining patterns.
-Specifically, we manually inspect and construct several verifiable constraint rules, in which the original instructions are appended with various constraints, and the responses are optionally modified to fulfill these constraints based on their original characteristics.
-Our predefined rules cover a wide range of diverse constraints ranging from high-level constraints, e.g., controlling the word frequency in the response, to lower-level constraints, e.g., setting specific wrapping formats [^1] The complete list of rules and descriptions can be found in the Appendix C.
-.
-Each rule contains (1) a set of constraint templates that serve as additional rule-instructions, on constraints the LLM should follow, and (2) specified operations that alternately edit the instruction and response to reach an alignment between them (i.e., all the constraints in the instruction hold in the response).
-For each sample from the original dataset, we randomly draw several rules to be applied in the editing.
-This produces an augmented sample with the constraints enforced that can be used for controllability tuning.
-An example is shown in Figure 1 (bottom).
-The original instruction is “What type of fruit would be a good addition to a fruit salad?”, and the corresponding response is “Grapes, kiwis, oranges, apples.”.
-If the constraint “Num of Nouns” is sampled and implemented on this sample, the augmented instruction becomes “What type of fruit would be a good addition to a fruit salad?
-Ensure the response has more than 3 nouns.”, and the original response remains unchanged on this type of rule.
-In this example, the constraint (number of nouns) is directly decided by the original data sample (4 nouns), thus avoiding the regeneration of response.
+(2024) in order to make them fulfilling multiple constraints, as shown in Figure 1 (top).
+However, the curation of new data is not only costly and inefficient, requiring careful editing by human experts or proprietary LLMs, but also represents a waste of previous efforts: It is impractical to discard all existing data and create brand new data every time we need to add more constraints to the instructions.
+Hence, we raise the question: Can we “recycle” existing SFT data without human/LLM editing and enforce various types of constraints in order to improve LLM controllability?
+Drawing inspiration from IFEval Zhou et al.
+(2023b), which utilizes verifiable constraints to evaluate LLMs’ controllability, and the human/model-free data augmentation in Mosaic-IT Li et al.
+(2024c), we propose Rule-based Data Recycling (RuleR), which automatically “re-engineering” existing SFT data for improving LLM controllability.
+As illustrated in Figure 1 (bottom), the key insight of RuleR is to automatically build constraint-augmented SFT dataset at no cost of human/LLM, by applying predefined rules to enforce constraints in the instructions and responses of the original dataset.
+Specifically, we manually inspect and construct a diverse set of rules as constraints, which specify the linguistic or formatting constraints on different parts of the response.
+To modify the original dataset, these rules are appended to the instructions, while the corresponding responses are modified if necessary to fulfill the constraints of these rules.
+Our predefined rules cover a wide range of diverse constraints generalizable to many application scenarios, ranging from high-level constraints, e.g., controlling the word frequency in the response, to lower-level constraints, e.g., setting specific wrapping formats of some keywords.
+Each rule is composed of (1) multiple templates to produce additional instructions enforcing the constraints, and (2) a piece of code [^1] https://github.com/tianyi-lab/RuleR that alternately edits the original instruction and response in order to make the edited response fulfill all the constraints appended to the instruction.
+For each sample from the original dataset, we randomly draw several rules to be applied to the editing.
+This produces an augmented sample with the constraints enforced so it can be used for controllability tuning.
+The complete list of rules and descriptions can be found in Appendix C and Table 4.
+Illustrative examples are provided in Figure 2, which showcase (a) a rule that constrains the number of letters of the response; (b) a rule that specifies the case of specific words (if presenting in the response); and (c) a rule that specifies the wrapping format of specific words.
+To ensure the consistency between the input constraints and the output response, we modify both the instruction and response based on the characteristics of the original data sample.
+For each sample, we only sample from the rules applicable to the original response, hence avoiding the potential incorrectness of the edited responses.
 Extensive experimental results on the IFEval (Instruction-Following Eval) Zhou et al.
-(2023b) benchmark on various datasets demonstrate the effectiveness of our human/model-free method in enhancing LLM controllability.
-Moreover, the instruction-following metrics also show that RuleR still preserves the general instruction-following ability promoted by the original SFT dataset.
-To the best of our knowledge, RuleR is the first human/model-free data augmentation approach designed to improve LLM controllability in enforcing multiple constraints to LLM-generated response.
-2 Method
+(2023b) benchmark on various base models and datasets demonstrate the effectiveness of RuleR in enhancing LLM controllability without extra help from human/model.
+On the other hand, RuleR still preserves the general instruction-following ability promoted by the original SFT dataset.
+This is demonstrated by the instruction-following metrics (Pair-wise comparison and Open LLM Leaderboard).
+To the best of our knowledge, RuleR is the first human/model-free data augmentation and recycling approach designed to improve LLM controllability under multiple constraints.
+2 Methodology
 2.1 Preliminaries
-Given a supervised finetuning dataset $D$, there are $n$ data samples each of which is represented by a tuple $(x_{i},y_{i})$, where $x_{i}$ represents the instruction and $y_{i}$ represents the corresponding response.
-Let $p_{\theta}(\cdot)$ denote the LLMs to be trained, with parameters $\theta$.
-In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by maximizing the following objective on each data $(x_{i},y_{i})$, $y_{i,j}$ represents the $j_{th}$ token of response $y_{i}$ and $l_{i}$ represents the token length:
-$\max_{\theta}\sum_{j=1}^{l_{i}}\log p_{\theta}\left(y_{i,j}|x_{i}\right),$ (1)
-2.2 Rule-based Data Recycling
-The basis for augmenting existing SFT datasets is to reconstruct instructions and responses.
-Most existing methods utilize human experts or strong teacher models for this process, which is costly, inefficient, and fails to get fully used to the existing data, as represented in Figure 1 (top).
-Inspired by Mosaic-IT which proposes a human/model-free augmentation method to reconstruct instruction-response pairs, we are motivated to design a rule-based method for controllability augmentation.
+Given a supervised finetuning dataset $D$, there are $n$ data samples, each represented by a tuple $(x_{i},y_{i})$, where $x_{i}$ represents the instruction and $y_{i}$ represents the corresponding response.
+Let $p_{\theta}(\cdot)$ denote the LLM with parameters $\theta$ to be trained.
+In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by maximizing the following objective on all the $N$ samples as $(x_{i},y_{i})$, in which $y_{i,j}$ represents the $j_{th}$ token of response $y_{i}$, $y_{i,<j}$ represents the tokens before $y_{i,j}$, and $l_{i}$ represents the token length of $y_{i}$:
+$\max_{\theta}\sum_{i=1}^{N}\sum_{j=1}^{l_{i}}\log p_{\theta}\left(y_{i,j}|x_{i},y_{i,<j}\right),$ (1)
+[Instruction]
+Ensure the 2nd paragraph is enclosed in "-| |-" in the response.
+Write a story of no more than 150 words about a knight who defends their kingdom from a dragon.
+Make sure the 2nd sentence is enclosed within "\$ \$", the 4th sentence is enclosed within "\$\$\$ \$", the 6th sentence is enclosed within "## #", the 8th sentence is enclosed within "*** *" in the response.
+[Response]
+In the kingdom of Arvania, there lived Sir Cedric, a valiant knight, sworn to defend the land against any threat.
+\$One day, the kingdom came under attack from a fire-breathing dragon, terrorizing the people and burning homes.\$ The king called on Sir Cedric, charging him to defeat the dragon and restore peace to the land.
+\$\$\$-|Wielding his sword and shield, Sir Cedric rode out to face the dragon, fearlessly engaging it in combat.\$ The battle was fierce, with the dragon’s flames clashing against the knight’s steel.
+##But Sir Cedric was determined and unrelenting, his bravery and skill unmatched.|- #
+Finally, with a mighty blow, Sir Cedric struck the dragon down, ending its reign of terror.
+***The people of Arvania rejoiced, celebrating the bravery of their hero.* And Sir Cedric returned in triumph, hailed as the dragon-slaying knight.
+[Instruction]
+Ensure the 1st sentence is in uppercase, the 2nd sentence is in uppercase, the 3rd sentence is in uppercase in the response.
+Keep the word "way" bounded by "@ @" in the response if it appears.
+Describe an example of a time you used to influence in a positive way
+Keep the 2nd sentence wrapped in "[[ ]]", the 4th sentence wrapped in "*** ***", and the 5th sentence wrapped in "# #" in the response.
+[Response]
+I RECENTLY HAD A TEAM PROJECT AT WORK WHERE I HAD TO INFLUENCE MY TEAM MEMBERS TO COMPLETE THE TASK IN THE MOST EFFICIENT @WAY@ POSSIBLE.
+[[I IDENTIFIED AREAS WHERE OUR TEAM COULD IMPROVE, AND I CONSTANTLY COMMUNICATED AND ENCOURAGED MY TEAMMATES TO STAY FOCUSED AND ON TRACK.]] I ALSO SOUGHT FEEDBACK ON HOW WE COULD IMPROVE THE PROCESS.
+***Through my efforts, I managed to motivate my team to complete the project ahead of schedule while also maintaining high-quality work.*** #This helped to boost morale and foster collaboration between our team members.# Everyone was very satisfied with the resulting outcome.
+Figure 3: Examples with multiple rules selected and implemented.
+The randomly generated rule-instructions are colored in violet.
+The upper example is augmented by $2$ different rules (Paragraph Wrapping, and Sentence Wrapping); the bottom example is augmented by $3$ different rules (Sentence Case, Keyword Wrapping, and Sentence Wrapping).
+The format differences in new responses are highlighted in red.
+2.2 Rule-based Data Recycling (RuleR)
+2.2.1 Rule Construction
+While most existing methods still require human experts or strong teacher LLMs to generate new data (Figure 1 (top)), we aim at “re-engineering” instructions and responses from existing SFT datasets for building controllability-focused datasets, without expensive and time-consuming supervision from humans or LLMs.
+In the following, we introduce a rule-based approach “Rule-based Data Recycling (RuleR)” to create high-quality data augmentations for improving LLM controllability.
+RuleR reformulates the original instructions and responses by applying rule-based edits according to pre-defined constraints.
+However, not every constraint can be applied to each sample without humans or LLMs.
+Hence, we only incorporate constraints compatible with the original sample and those can be implemented with simple rectifications like regular expressions.
+In RuleR, the rules and constraints are elaborately designed and selected to specify or change certain characteristics of the original data, resulting in consistency between each constraint-augmented instruction and the reformulated response.
+Specifically, we focus on the characteristics of responses that can be defined by rules, e.g., by checking the 220 distinct linguistic features in the LFTK package Lee and Lee (2023).
+In addition, we collect constraints from existing works and widely used instruction-tuning datasets.
+These diverse characteristics include punctuation-, word-, sentence-, paragraph-level occurrences, and frequencies.
+We construct rules constraining or specifying the characteristics of the original response.
+In conjunction with the rules containing these characteristics, we also create rules specifying the format of responses to improve LLM’s format-following capability.
+To ensure that the rules selected or sampled for each sample are applicable and legal, we apply the following additional protocols: (1) The rules need to be applicable to the original sample.
+For instance, the rule “Generating a title before giving the response” is not applicable as we can not generate a title without the help of humans or other additional models.
+(2) The rules should not remove the content of the original response.
+For instance, the rule “Ensure the word xxx is not shown in the response” is not applicable since we can not directly remove this word from the response as it is part of the original response.
+(3) The rules should be compatible with the original sample.
+If the rule is “Ensure there are more than N sentences in the response”, then it can not be applied to samples whose responses have $<N$ sentences.
+The complete list of rules is provided in Table 4, which covers both high-level constraints such as the term frequency, and lower-level constraints such as specific wrapping formats.
+2.2.2 Rule Implementation
+To implement each rule as a program editing instructions and responses, we formulate each pre-defined rule as a tuple, $(\mathbf{S}_{k},f_{k},g_{k})$, where $\mathbf{S}_{k}$ represents the set of manually curated templates for creating instructions of the $k^{th}$ rule, while $f_{k}$ and $g_{k}$ are the corresponding functions to reformulate instructions and responses (if necessary), respectively.
+Specifically, the function $f_{k}$ selects an appropriate template of a rule for a given data sample and augments the original instruction with an instruction generated by the template.
+It first selects a subset of rules applicable to the characteristics of the response (e.g., presence of keywords, number of sentences, etc.).
+Then, we randomly draw one rule out of the subset and create a formatted instruction of the rule from the template.
+Such rule instruction is appended to the original instruction.
+Specifically, for each data sample $(x_{i},y_{i})$, the augmented instruction $x_{i,aug}$ is reformulated according to the characteristics of the original sample and the corresponding template sets, i.e.,
+$x_{i,aug}=f_{k}(x_{i},y_{i},\mathbf{S}_{k}),$ (2)
+The function $g_{k}$ is designed to modify the response to be consistent with the augmented instruction (after applying function $f_{k}$), i.e., with the rule applied.
+Applying $g_{k}$ either preserves the original response or revises it.
+Some rules do not require editing of responses, e.g., keyword appearance, the number of nouns, etc.
+For rules defining the case or format of certain parts in the response, modifications are needed.
+The detailed descriptions of each predefined rule can be found in Appendix C.
+Specifically, the augmented response $y_{i,aug}$ will be optionally modified based on the selected rule $k$:
+$y_{i,aug}=g_{k}(x_{i},y_{i},\mathbf{S}_{k}).$ (3)
+With the augmented sample $(x_{i,aug},y_{i,aug})$, the training objective becomes:
+$\max_{\theta}\sum_{i=1}^{N}\sum_{j=1}^{l_{i,aug}}\log p_{\theta}\left(y_{i,aug,j}|x_{i,aug},y_{i,aug,<j}\right),$ (4)
+where $y_{i,aug,j}$ represents the $j$th token of response $y_{i,aug}$ and $l_{i,aug}$ represents its token length.
+2.2.3 Multi-rule Implementation
+§ 2.2.2 illustrates how RuleR applies one rule to edit each data sample.
+To create more complex, diverse, and challenging samples, we can extend it to multiple rules randomly drawn from the feasible set of rules.
+We provide examples of multi-rule augmentation in Figure 3.
+It shows that the complexity of the augmented data grows dramatically, forcing the LLMs to learn to follow multiple constraints.
 Model Dataset Method Instruction Following Eval Pair-wise Open LLM Leaderboard
@@
 DEITA 6K Baseline 41.22 51.08 44.55 54.92 1.000 64.82 60.41 82.52 61.57 54.76
-Ours 45.10 54.32 47.69 56.95 1.010 65.34 61.16 82.78 62.74 54.66
+Ours 42.14 52.28 46.77 56.59 1.010 65.43 61.86 82.71 62.66 54.49
 Llama2-7B Alpaca-GPT4 Baseline 26.25 36.33 30.31 40.29 1.000 58.71 54.69 80.05 47.89 52.21
@@
 Ours 45.47 58.15 50.09 61.99 1.010 60.84 58.28 82.37 54.35 48.36
-Table 1: The comparisons on the Instruction Following Eval, Pair-wise Comparison Winning Score, and the Open LLM Leaderboard, on 3 different base models and several different instruction tuning datasets.
-The baselines are models trained with the original dataset.
-Prompt and Inst represent Prompt-level and Instruction-level accuracy; S and L represent Strict and Loose versions.
-A, H, M, and T represent ARC, HellaSwag, MMLU, and TruthfulQA.
-With this goal, the first challenge is the reformulation of instructions.
-Due to the lack of human experts or LLMs to generate the corresponding responses for any given instructions, it is not feasible to add arbitrary constraints to the original instructions.
-To solve this problem, we propose to only incorporate constraints that are compatible with the original data sample.
-Specifically, we propose to enhance LLM controllability by Rule-based Data Recycling (RuleR), in which the rules and constraints used to reformulate instruction are determined by the original data, ensuring the alignment between instruction and response.
-Motivated by IFEval, we manually inspect and propose several verifiable rules in which the responses are not required for regeneration by LLMs or human experts, but only need optional slight adjustments.
-The complete list of rules is listed in Table 3 ranging from high-level constraints like term frequency to lower-level constraints like specific wrapping formats.
-During implementation, we instantiate each predefined rule as a tuple , $(\mathbf{S}_{k},f_{k},g_{k})$, where $\mathbf{S}_{k}$ represents the set of manually curated rule-instruction templates for the $k^{th}$ rule, $f_{k}$ and $g_{k}$ represent the function to reformulate instructions and optionally modify responses respectively, as shown in Figure 1(bottom).
-For each data sample $(x_{i},y_{i})$, the augmented instruction $x_{i,aug}$ will be reformulated according to the characteristics of the original data and the corresponding template sets:
-$x_{i,aug}=f_{k}(x_{i},y_{i},\mathbf{S}_{k})$
-where specifically, a random rule-instruction template will be sampled and filled with the corresponding characteristics as the additional constraint to the original instruction.
-Then the rule-instruction is concatenated with the original instruction becoming the augmented instruction.
-Similarly, the augmented response $y_{i,aug}$ will be optionally modified based on the selected rule $k$:
-$y_{i,aug}=g_{k}(x_{i},y_{i},\mathbf{S}_{k}),$
-where the function $g_{k}$ varies dramatically for different rules: For some of the rules, no modification needs to be done to the original response, while for others, some arbitrary part of the response needs to be replaced with the desired user format.
-3 Experiments
-3.1 Experimental Setup
-We conduct experiments on several datasets including the Alpaca Taori et al.
-(2023), Alpaca-GPT4 Peng et al.
-(2023), WizardLM Xu et al.
-(2023), Recycled WizardLM Li et al.
-(2023a), DEITA Liu et al.
-(2023), and on several base LLMs including Llama2-7B, Llama2-13B Touvron et al.
-(2023b), and Mistral-7B Jiang et al.
-(2023), introduction on datasets can be found in the Appendix A.
-We utilize the prompt and code base from Vicuna Chiang et al.
-(2023a) and flash attention Dao et al.
-(2022).
+Table 1: Main Results.
+Evaluation on the Instruction Following Eval, Pair-wise Comparison Winning Score, and the Open LLM Leaderboard.
+We compare RuleR with Baseline for finetuning three base models on several different instruction tuning datasets.
+Baseline – models trained with the original dataset; Ours – models trained with RuleR-recycled datasets; Prompt – Prompt-level accuracy; Inst – Instruction-level accuracy; S and L represent Strict and Loose versions.
+A, H, M, and T denote ARC, HellaSwag, MMLU, and TruthfulQA.
+3 Experimental Setup
+3.1 Implementation Details
+We utilize the prompt and code base from Vicuna Chiang et al. (2023a) and flash attention Dao et al. (2022) for all our experiments.
 The Adam optimizer Kingma and Ba (2017) is utilized with the batch size of $128$ and with the max token length of $2048$.
@@
 All other configurations are kept the same as the baselines.
-3.2 Experimental Results
+3.2 Training Datasets
+We utilize 5 SFT datasets to evaluate the effectiveness of our method: Alpaca dataset Taori et al.
+(2023), Alpaca-GPT4 dataset Peng et al.
+(2023), WizardLM dataset Xu et al.
+(2023), Recycled WizardLM Dataset Li et al.
+(2023a), DEITA dataset Liu et al.
+(2023).
+A detailed illustration of datasets can be found in the Appendix A.
+3.3 Evaluation Metrics
+We employ 3 commonly accepted metrics for the evaluation, including IFEval (Instruction-Following Eval), Pair-wise Comparison, and Open LLM Leaderboard.
+IFEval Zhou et al.
+(2023b) is the primary evaluation metric employed in our study due to its compatibility with our motivation.
+It focuses on evaluating how LLMs follow various additional constraints, such as specifying a word count or requiring the inclusion of certain keywords a specified number of times.
+To avoid the utilization of LLMs during evaluation, it proposes 25 distinct types of verifiable instructions.
+There are 541 prompts in total and each of them incorporates one or more of these verifiable instructions, ensuring the comprehensiveness of the evaluation.
+IFEval serves as a great inspiration for our method, and there exist semantical overlappings between their verifiable instructions and our rules.
+However, the specific prompts used in IFEval are kept unknown in the construction of our rule templates, avoiding potential template leakage.
+Moreover, IFEval only needs to verify the responses, while our method needs to modify the responses for the training, which pushes this process a step further.
+Consequently, this benchmark not only facilitates a comprehensive comparison but also provides valuable insights that align with our purpose.
+Pair-wise Comparison involves evaluating responses from LLMs like GPT-4, especially in open-domain contexts.
+This method has shown a notable alignment with human assessments, providing a credible evaluative foundation Zheng et al.
+(2024); Li et al.
+(2023b); Sottana et al.
+(2023).
+We utilize test instruction sets from WizardLM Xu et al.
+(2023) and Vicuna Chiang et al.
+(2023b), comprising 218 and 80 diverse human-curated instructions, respectively.
+Following the framework by Li et al.
+(2024e, d), we prompt judging LLM to rate responses on a scale from 1 to 10 across multiple dimensions such as relevance and accuracy.
+To mitigate positional bias Ko et al.
+(2020); Wang et al.
+(2023a), responses are presented in two distinct sequences.
+Evaluation outcomes are categorized into the ’win-tie-lose’ system for each instruction.
+One model’s response can be noted as Win/Lose only if it has a Higher/Lower score in both directions.
+The specific evaluation prompt is presented in the Appendix B.
+Open LLM Leaderboard employs Eval Harness Gao et al.
+(2021a) to systematically assess generative language models through diverse tasks.
+Key benchmarks include ARC Clark et al.
+(2018), HellaSwag Zellers et al.
+(2019), MMLU Hendrycks et al.
+(2020), and TruthfulQA Lin et al.
+(2021), collectively evaluating reasoning abilities, common-sense knowledge, and factual accuracy.
+This approach highlights LLM strengths and areas for improvement in critical aspects of language understanding and generation Gao et al.
+(2021b).
+4 Experimental Results
+4.1 Main Results
 The main experimental results are presented in Table 1, containing the performance comparison on the Instruction Following Eval Zhou et al.
@@
 The Pair-wise comparison and Open LLM leaderboard results showcase comparable or sometimes better performances compared with the baseline models.
-We hypothesize that the additional constraints might complicate the original instructions, thus leading to potentially improved instruction-following abilities.
+We hypothesize that the additional constraints largely complicate the original instructions, as shown in Figure 3, thus forcing the LLMs to understand each constraint before generating responses, thus leading to potentially improved instruction-following abilities.
 4.2 Ablation Studies
@@
 Aug Rate = 0.9 (*) 47.46 1.044 61.05
-Table 2: Ablations.
-The experiments are conducted on Mistral-7B on Alpaca-GPT4 data and the configuration notated with “(*)” represents our default setting.
-“IF Eval” and “Open LLM” represent the averaged scores.
+Table 2: Ablation Study. Average scores on “IF Eval” and “Open LLM” of Mistral-7B finetuned with ours-augmented Alpaca-GPT4 data. “(*)” represents default configuration.
+Sub-Category Case Combination Content Json Keywords Language Length Punctuation Start End
+Baseline (Strict) 22.47 16.92 75.47 63.06 44.79 58.06 31.47 12.12 58.21
+Aug Rate = 0.1 78.65 63.08 45.28 59.24 33.13 74.19 32.17 30.30 79.10
+Aug Rate = 0.3 76.40 58.46 45.28 45.86 30.06 58.06 35.66 22.73 71.64
+Aug Rate = 0.5 70.79 69.23 49.06 45.22 39.88 77.42 28.67 13.64 61.19
+Aug Rate = 0.7 74.16 61.54 43.40 47.77 33.13 51.61 30.77 77.27 79.10
+Aug Rate = 0.9 67.42 75.38 47.17 47.13 34.97 61.29 31.47 66.67 82.09
+Baseline (Loose) 24.72 20.00 75.47 66.88 48.47 64.52 37.06 15.15 58.21
+Aug Rate = 0.1 79.78 69.23 45.28 61.15 37.42 74.19 36.36 40.91 80.60
+Aug Rate = 0.3 77.53 66.15 45.28 47.13 36.81 64.52 39.16 25.76 76.12
+Aug Rate = 0.5 77.53 70.77 49.06 45.86 42.94 80.65 32.17 15.15 65.67
+Aug Rate = 0.7 77.53 70.77 43.40 48.41 35.58 58.06 35.66 84.85 80.60
+Aug Rate = 0.9 70.79 80.00 47.17 48.41 40.49 67.74 39.16 69.70 85.07
+Table 3: Sub-category performance on IFEval benchmark of Mistral-7B finetuned with RuleR-augmented Alpaca-GPT4 data.
+The top section represents the performance by the strict criterion while the bottom represents the loose.
 In this section, ablation experiments are conducted on Mistral-7B with the Alpaca-GPT4 dataset.
-As shown in Table 2, “Single Temp.” represents utilizing only one rule-insturtcion template for each rule, while “Diverse Temp.” represents utilizing several different templates (with the same meaning) for each rule and those templates are randomly sampled during the augmentation.
-The superiority of “Diverse Temp.” showcases the necessity of providing diverse templates to avoid potential overfitting.
+The experiments aim to evaluate the impact of several factors on our method: template diversity, the maximum number of rules utilized in each data sample, and the probability of applying augmentation to each sample.
+The performance of each configuration is evaluated on three evaluation metrics, Instruction Following Eval Pair-wise winning scores, and Open LLM Leaderboard, to simultaneously inspect the constraint controllability and general instruction-following ability.
+4.2.1 Effect of Templates
+This section illustrates the effectiveness of utilizing diverse rule templates for the augmentation.
+As shown in Table 2, “Single Temp.” represents utilizing only one rule-instruction template for each rule, while “Diverse Temp.” represents utilizing several different templates, approximately $10$, with the same meaning for each rule and those templates are randomly sampled during the augmentation.
+As indicated in the table, “Diverse Temp.” demonstrates superior performance in both IF Eval and Pair-wise winning scores, with slightly lower accuracy on the Open LLM leaderboard.
+This result suggests that “Diverse Temp.” enhances the model’s constraint controllability while enhancing its general instruction-following capabilities compared to “Single Temp.” On the contrary, when fixing the rule templates to one single template, potential overfitting to the template might occur and thus negatively influence the model performances.
+4.2.2 Effect of Rule Numbers
+This section illustrates the effects of implementing multiple random rules together in one single sample.
 “Max Rule = $x$” represents the setting in which at most $x$ different rules can be sampled and utilized on each original data sample.
-The performances of models, whose Max Rule is set under 5, are promising, while the performance declines on all 3 evaluation metrics when the Max Rule is set as 5.
-These results indicate that too many rules utilized together in one sample might be harmful to LLM capability.
-“Aug Rate = x” represents there is a probability of x to apply our augmentation to each sample.
-It can be observed that increasing the augmentation rate is beneficial to the controllability of LLMs, while slightly harmful to the general instruction-following capability.
+In the augmentation process, a random value will be sampled in the range of $[0,x]$ as the number of rules in this sample.
+The examples in Figure 2 showcase the scenario when only one rule is implemented and examples in Figure 3 showcase the scenario when multiple rules are implemented.
+Compared to the baseline, nearly all settings show performance improvements across the three evaluation metrics.
+However, as shown in the table, the IF Eval score initially increases, reaching its peak when “Max Rule = $3$”, before declining.
+The Pair-wise score, on the other hand, consistently decreases from “Max Rule = $1$” to “Max Rule = $5$”.
+Although no clear trend is observed on the Open LLM leaderboard, Max Rule = 5 shows the worst performance.
+These results suggest that applying too many rules to a single sample may impair the LLM’s capability, even though sampling and applying multiple rules can be beneficial when done in moderation, which might be because the original instruction becomes so complex for LLMs to learn.
+4.2.3 Effect of Augmentation Rate
+This section illustrates the effect of the augmentation rate in our method. “Aug Rate = $x$” represents there is a probability of $x$ to apply our augmentation to each sample.
+It is observed that as the augmentation rate increases, performance improves on the IF Eval and mostly improves on the Pair-wise evaluation.
+However, the results on the Open LLM leaderboard show limited improvement from the baseline, without a significant upward trend.
+This phenomenon indicates that increasing the augmentation rate primarily enhances the LLM’s constraint controllability, and it also has a positive impact on its general instruction-following capability.
 However, the gaps on IF Eval are much larger than the other 2 metrics, indicating this rate will mostly influence the multi-constraint controllability of LLM.
+4.3 Detailed Sub-Category Analysis
+In this section, detailed comparisons between our models and baseline models are provided in Table 3 for further analysis of the effects of our method.
+It is worth noting that the specific prompts used in IFEval are kept unknown in the construction of our rule templates, avoiding potential overfitting to the templates.
+Models trained with our recycled data outperform the baseline model consistently in “Case”, “Combination”, “Punctuation”, and “Start End” categories, which are partially well-recycled by our method.
+In the “Length” category, our models are only slightly better than the baseline model although our recycling method contains this kind of constraints.
+After further investigation, we find the performance in this category is mainly influenced by the original data length distributions.
+Since our method does not introduce more data samples, thus not able to improve the performance dramatically, but the better performance indeed provides the model with a better understanding of response length.
+Interestingly, the performance in the “Language” category also shows consistent improvement although we do not introduce any more data.
+Considering the consistency between this performance and the Pair-wise comparison performance, we hypothesize this improvement is caused by the better general instruction-following abilities provided by our method.
+The performance in the “Content” category presents one of the limitations of our Rule-based Recycling method, without utilizing other models or human experts to rewrite the instruction and response, it’s hard for our method to modify the content of the existing response.
+The performances of our models in the “Json” and “Keywords” categories are merely slightly lower, which is mainly affected by the diversity of original training datasets.
+Comparing the performance changes across augmentation rates, LLMs obtain better performances when the augmentation rate is higher, except for “Case”, indicating the easiness of case-related constraints for LLMs to understand and learn.
 5 Conclusion
-In this work, we proposed a novel method, Rule-based Data Recycling (RuleR) which utilizes the constraints derived from the existing data to enhance the dataset and incorporates these constraints into the training process, thereby improving the multi-constraint controllability of LLMs.
-This approach presents a promising direction for developing SFT-augmenting methods to improve LLM controllability, paving the way for more efficient advancements of LLM.
+In this work, we proposed Rule-based Data Recycling (RuleR), which modified the original instructions and responses from an existing dataset by rule-defined constraints.
+The “recycled” data aims to enhance the LLMs’ capability to generate outputs fulfilling the constraints specified in the input, thereby improving the controllability of LLMs.
+RuleR took the first step of exploring rule-based data recycling.
+As a data generation approach, it is much cheaper than existing methods generating new data from scratch since RuleR does not rely on any human supervision or additional models.
+Our RuleR can serve as a plug-and-play and easy-to-use method that converts any existing instruction-tuning datasets to new datasets for gaining better controllability.
+Hence, our method is more accessible and widely applicable.
+RuleR provides a promising and sustainable strategy to continuously improve LLMs.
+Moreover, our method has great potential to be extended to other settings.
+For example, with the help of additional models, we can apply content-level or semantic-level rules to produce more diverse training data, as discussed in the next section.
 Limitations
-Our method focuses on improving LLM controllability by rule-based data augmentation to avoid extra regeneration processes from humans or models.
-Though the alleviation of reliance on human/model editing is our main contribution, it inevitably limits the types of constraints that can be implemented on the original data sample.
-Currently, all the constraint rules are based on verifiable shallow syntactic characteristics like the occurrence and frequency of words or sentences, etc, lacking constraint rules that focus more on the semantic feature or content.
-Thus, our method shows great potential to be further enhanced if further involves small models, which will still remain efficient and avoid the regeneration process.
+Our method focuses on improving LLM controllability by rule-based editing of existing data, thereby avoiding the extra cost of data generation by humans or expert models.
+Though it saves the cost of human/model editing, the rules inevitably limit the types of constraints that can be applied to modify the original data.
+In the presented RuleR, all the constraints and rules are based on verifiable shallow syntactic characteristics such as the occurrence and frequency of words or sentences while lacking constraints and controllability on the semantic feature or content.
+This implies a potential of RuleR to be further enhanced by modifying the semantic content with a smaller model, which retains a comparable efficiency of rule-based editing.
 Appendix A Training Dataset
@@
 Appendix B Evaluation Metrics
-We employ 3 commonly accepted metrics for the evaluation, including IFEval (Instruction-Following Eval), Pair-wise Comparison, and Open LLM Leaderboard
-IFEval Zhou et al.
-(2023b) is the primary evaluation metric employed in our study due to its compatibility with our motivation.
-It focuses on evaluating how LLMs follow various additional constraints, such as specifying a word count or requiring the inclusion of certain keywords a specified number of times.
-To avoid the utilization of LLMs during evaluation, it proposes 25 distinct types of verifiable instructions.
-There are 541 prompts in total and each of them incorporates one or more of these verifiable instructions, ensuring the comprehensiveness of the evaluation.
-IFEval serves as a great inspiration for our method, and there exist semantical overlappings between their verifiable instructions and our rules.
-However, the specific prompts used in IFEval are kept unknown in the construction of our rule templates, avoiding potential data leakage.
-Moreover, IFEval only needs to verify the responses, while our method needs to modify the responses for the training, which pushes this process a step further.
-Consequently, this benchmark not only facilitates a comprehensive comparison but also provides valuable insights that align with our purpose.
-Pair-wise Comparison involves evaluating responses from LLMs like GPT-4, especially in open-domain contexts.
-This method has shown a notable alignment with human assessments, providing a credible evaluative foundation Zheng et al.
-(2024); Li et al.
-(2023b); Sottana et al.
-(2023).
-We utilize test instruction sets from WizardLM Xu et al.
-(2023) and Vicuna Chiang et al.
-(2023b), comprising 218 and 80 diverse human-curated instructions, respectively.
-Following the framework by Li et al.
-(2024e, d), we prompt judging LLM to rate responses on a scale from 1 to 10 across multiple dimensions such as relevance and accuracy.
-To mitigate positional bias Ko et al.
-(2020); Wang et al.
-(2023a), responses are presented in two distinct sequences.
-Evaluation outcomes are categorized into the ’win-tie-lose’ system for each instruction.
-One model’s response can be noted as Win/Lose only if it has a Higher/Lower score in both directions.
-The specific evaluation prompt is presented in Figure 2.
-Open LLM Leaderboard employs Eval Harness Gao et al.
-(2021a) to systematically assess generative language models through diverse tasks.
-Key benchmarks include ARC Clark et al.
-(2018), HellaSwag Zellers et al.
-(2019), MMLU Hendrycks et al.
-(2020), and TruthfulQA Lin et al.
-(2021), collectively evaluating reasoning abilities, common-sense knowledge, and factual accuracy.
-This approach highlights LLM strengths and areas for improvement in critical aspects of language understanding and generation Gao et al.
-(2021b).
+The prompt for pair-wise comparison is shown in Figure 4.
   Prompt for Performance Evaluation
@@
 Appendix C Predefined Rules
-Rule Type Rule Name Example Rule-Instruction Templates
+Rule Type Rule Name Example of an Instruction Template for the Rule
 Keyword Frequency Keyword Appearance Ensure {Keyword} is in the response.
@@
 Table 4: The list of predefined constraint rules.
+Each rule contains (1) a set of constraint templates that serve as additional rule-instructions, on constraints the LLM should follow, and (2) specified methods that alternately edit the instruction and response to reach an alignment between them.
 In this section, we will dive into the predefined rules describing each constraint specifically.

### v4

@@
 Despite the remarkable advancement of Large language models (LLMs), they still lack delicate controllability under sophisticated constraints, which is critical to enhancing their response quality and the user experience.
-While conditional supervised fine-tuning (SFT) can potentially improve LLM controllability, curating new SFT data to fulfill the constraints usually relies on human experts or proprietary LLMs, which is time-consuming and expensive.
+While supervised fine-tuning (SFT) can potentially improve LLM controllability, curating new SFT data to fulfill the constraints usually relies on human experts or proprietary LLMs, which is time-consuming and expensive.
 To bridge this gap, we propose Rule-based Data Recycling (RuleR), a human/LLM-free data augmentation method incorporating multiple constraints into the original SFT data.
 Instead of creating new responses from scratch, RuleR integrates linguistic or formatting rules into the original instructions and modifies the responses to fulfill the rule-defined constraints.
-Training on the “recycled” data consolidates LLM capability to generate constrained outputs.
-Extensive experiments demonstrate RuleR’s effectiveness in improving LLM controllability while maintaining general instruction-following performance.
-RuleR’s code is released on https://github.com/tianyi-lab/RuleR.
+Training on the “recycled” data consolidates LLM capability to generate constrained outputs, improving LLM controllability while maintaining promising general instruction-following capabilities.
 RuleR: Improving LLM Controllability by Rule-based Data Recycling
-Ming Li*1, Han Chenthanks: Equal Contribution., Chenguang Wang*2, Dang Nguyen1, Dianqi Li, Tianyi Zhou1 1University of Maryland 2Stony Brook University {minglii, tianyi}@umd.edu
+Ming Li*1, Han Chenthanks: Equal Contribution., Chenguang Wang*2, Dang Nguyen1, Dianqi Li, Tianyi Zhou1 1University of Maryland 2Stony Brook University {minglii, tianyi}@umd.edu Project: https://github.com/tianyi-lab/RuleR
 1 Introduction
 Figure 1: Comparing widely-used data generation strategy (top) and RuleR (bottom) enhancing LLM controllability.
-Most existing methods rely on human/model supervision or reflection to generate new instructions and responses.
-However, discarding existing data is a waste of efforts taken to curate them by human experts or expert LLMs.
-On the contrary, our RuleR demonstrates that simple rule-based (human/model-free) editing of existing data can greatly improve LLM controllability.
-Figure 2: Examples of the data recycling workflows in Rule-based Data Recycling (RuleR).
-(a), (b) and (c) select different predefined rules to modify the original data to fulfill constraints on the complexity or format of the response.
-The differences in new responses are highlighted in red, the example in (a) has already satisfied the appended constraint, thus the response is kept unchanged.
-Despite the remarkable advancement of the current Large language models (LLMs) and the continuous efforts to build high-quality supervised fine-tuning (SFT) datasets Brown et al.
-(2020); OpenAI (2023); Touvron et al.
-(2023a, b); Jiang et al.
-(2023), one critical challenge is to generate responses better interacting with humans, with the utility and effectiveness maximized for end-users Liu et al.
-(2024).
+Most existing methods rely on human/model rewriting to generate new instructions and responses.
+However, discarding existing data is a waste of effort.
+Our RuleR demonstrates that simple rule-based (human/model-free) editing of existing data can generate new SFT data that improves LLM controllability.
+Despite the remarkable advancement of the current Large language models (LLMs) and the continuous efforts to build high-quality supervised fine-tuning (SFT) datasets, one critical challenge is to generate responses better interacting with humans, with the utility and effectiveness maximized for end-users Liu et al.
+(2024); Huang et al.
+(2024, 2023).
 According to the systematic investigation from Liu et al.
 (2024), it is essential for LLMs to constrain their outputs to follow user-specified formats or characteristics.
-In various practical applications, without any constraint or format being enforced, free-formed responses are not legal or directly applicable.
+In various practical applications, free-formed responses are not legal or directly applicable without any constraint or format being enforced.
 It has also been verified on LLM Agents Li et al.
@@
 (2023, 2024b); Zhang et al.
-(2024) that enforcing predefined formats is necessary for various tasks.
+(2024) that enforcing predefined formats is necessary for tasks.
 However, existing SFT datasets are mainly composed of general instructions without user-specified constraints Wei et al.
@@
 (2023a, 2024b, 2024a); He et al.
+(2024); Dong et al.
+(2024); Wu et al.
 (2024) in order to make them fulfilling multiple constraints, as shown in Figure 1 (top).
@@
 (2023b), which utilizes verifiable constraints to evaluate LLMs’ controllability, and the human/model-free data augmentation in Mosaic-IT Li et al.
-(2024c), we propose Rule-based Data Recycling (RuleR), which automatically “re-engineering” existing SFT data for improving LLM controllability.
-As illustrated in Figure 1 (bottom), the key insight of RuleR is to automatically build constraint-augmented SFT dataset at no cost of human/LLM, by applying predefined rules to enforce constraints in the instructions and responses of the original dataset.
+(2024c), we propose Rule-based Data Recycling (RuleR), which automatically “recycles” existing SFT data for improving LLM controllability.
+As illustrated in Figure 1 (bottom), the key insight of RuleR is to automatically build constraint-augmented SFT datasets at no cost of human/LLM efforts, by applying predefined rules to the original instructions and responses.
 Specifically, we manually inspect and construct a diverse set of rules as constraints, which specify the linguistic or formatting constraints on different parts of the response.
-To modify the original dataset, these rules are appended to the instructions, while the corresponding responses are modified if necessary to fulfill the constraints of these rules.
 Our predefined rules cover a wide range of diverse constraints generalizable to many application scenarios, ranging from high-level constraints, e.g., controlling the word frequency in the response, to lower-level constraints, e.g., setting specific wrapping formats of some keywords.
-Each rule is composed of (1) multiple templates to produce additional instructions enforcing the constraints, and (2) a piece of code [^1] https://github.com/tianyi-lab/RuleR that alternately edits the original instruction and response in order to make the edited response fulfill all the constraints appended to the instruction.
+Each rule is composed of (1) multiple templates to produce additional instructions enforcing the constraints, and (2) a piece of code that alternately edits the original instruction and response in order to make the edited response fulfill all the constraints appended to the instruction.
 For each sample from the original dataset, we randomly draw several rules to be applied to the editing.
 This produces an augmented sample with the constraints enforced so it can be used for controllability tuning.
-The complete list of rules and descriptions can be found in Appendix C and Table 4.
-Illustrative examples are provided in Figure 2, which showcase (a) a rule that constrains the number of letters of the response; (b) a rule that specifies the case of specific words (if presenting in the response); and (c) a rule that specifies the wrapping format of specific words.
+The complete list of rules and descriptions can be found in Appendix E.
+Illustrative examples are provided in Figure 2, which showcase (a) a rule that constrains the number of letters; (b) a rule that specifies the case of specific words (if presenting in the response); and (c) a rule that specifies the wrapping format of specific words.
 To ensure the consistency between the input constraints and the output response, we modify both the instruction and response based on the characteristics of the original data sample.
 For each sample, we only sample from the rules applicable to the original response, hence avoiding the potential incorrectness of the edited responses.
-Extensive experimental results on the IFEval (Instruction-Following Eval) Zhou et al.
-(2023b) benchmark on various base models and datasets demonstrate the effectiveness of RuleR in enhancing LLM controllability without extra help from human/model.
+Figure 2: Examples of our data recycling workflows.
+(a), (b) and (c) select different predefined rules to modify the original data to fulfill constraints on the complexity or format of the response.
+The differences in new responses are highlighted in red, the example in (a) has already satisfied the appended constraint, thus the response is kept unchanged.
+Extensive experimental results on the IFEval benchmark on various base models and datasets demonstrate the effectiveness of RuleR in enhancing LLM controllability without extra help from humans/models.
 On the other hand, RuleR still preserves the general instruction-following ability promoted by the original SFT dataset.
@@
 $\max_{\theta}\sum_{i=1}^{N}\sum_{j=1}^{l_{i}}\log p_{\theta}\left(y_{i,j}|x_{i},y_{i,<j}\right),$ (1)
-[Instruction]
-Ensure the 2nd paragraph is enclosed in "-| |-" in the response.
-Write a story of no more than 150 words about a knight who defends their kingdom from a dragon.
-Make sure the 2nd sentence is enclosed within "\$ \$", the 4th sentence is enclosed within "\$\$\$ \$", the 6th sentence is enclosed within "## #", the 8th sentence is enclosed within "*** *" in the response.
-[Response]
-In the kingdom of Arvania, there lived Sir Cedric, a valiant knight, sworn to defend the land against any threat.
-\$One day, the kingdom came under attack from a fire-breathing dragon, terrorizing the people and burning homes.\$ The king called on Sir Cedric, charging him to defeat the dragon and restore peace to the land.
-\$\$\$-|Wielding his sword and shield, Sir Cedric rode out to face the dragon, fearlessly engaging it in combat.\$ The battle was fierce, with the dragon’s flames clashing against the knight’s steel.
-##But Sir Cedric was determined and unrelenting, his bravery and skill unmatched.|- #
-Finally, with a mighty blow, Sir Cedric struck the dragon down, ending its reign of terror.
-***The people of Arvania rejoiced, celebrating the bravery of their hero.* And Sir Cedric returned in triumph, hailed as the dragon-slaying knight.
-[Instruction]
-Ensure the 1st sentence is in uppercase, the 2nd sentence is in uppercase, the 3rd sentence is in uppercase in the response.
-Keep the word "way" bounded by "@ @" in the response if it appears.
-Describe an example of a time you used to influence in a positive way
-Keep the 2nd sentence wrapped in "[[ ]]", the 4th sentence wrapped in "*** ***", and the 5th sentence wrapped in "# #" in the response.
-[Response]
-I RECENTLY HAD A TEAM PROJECT AT WORK WHERE I HAD TO INFLUENCE MY TEAM MEMBERS TO COMPLETE THE TASK IN THE MOST EFFICIENT @WAY@ POSSIBLE.
-[[I IDENTIFIED AREAS WHERE OUR TEAM COULD IMPROVE, AND I CONSTANTLY COMMUNICATED AND ENCOURAGED MY TEAMMATES TO STAY FOCUSED AND ON TRACK.]] I ALSO SOUGHT FEEDBACK ON HOW WE COULD IMPROVE THE PROCESS.
-***Through my efforts, I managed to motivate my team to complete the project ahead of schedule while also maintaining high-quality work.*** #This helped to boost morale and foster collaboration between our team members.# Everyone was very satisfied with the resulting outcome.
-Figure 3: Examples with multiple rules selected and implemented.
-The randomly generated rule-instructions are colored in violet.
-The upper example is augmented by $2$ different rules (Paragraph Wrapping, and Sentence Wrapping); the bottom example is augmented by $3$ different rules (Sentence Case, Keyword Wrapping, and Sentence Wrapping).
-The format differences in new responses are highlighted in red.
 2.2 Rule-based Data Recycling (RuleR)
 2.2.1 Rule Construction
-While most existing methods still require human experts or strong teacher LLMs to generate new data (Figure 1 (top)), we aim at “re-engineering” instructions and responses from existing SFT datasets for building controllability-focused datasets, without expensive and time-consuming supervision from humans or LLMs.
-In the following, we introduce a rule-based approach “Rule-based Data Recycling (RuleR)” to create high-quality data augmentations for improving LLM controllability.
+While most existing methods still require human experts or strong teacher LLMs to generate new data (Figure 1 (top)), we aim at “recycling” instructions and responses from existing SFT datasets to build controllability-focused datasets, without expensive and time-consuming supervision from humans or LLMs.
+In the following, we introduce a rule-based approach “Rule-based Data Recycling (RuleR)” to create high-quality augmented data for improving LLM controllability.
 RuleR reformulates the original instructions and responses by applying rule-based edits according to pre-defined constraints.
-However, not every constraint can be applied to each sample without humans or LLMs.
+However, not every constraint is applicable to a randomly selected sample without fully rewriting.
 Hence, we only incorporate constraints compatible with the original sample and those can be implemented with simple rectifications like regular expressions.
-In RuleR, the rules and constraints are elaborately designed and selected to specify or change certain characteristics of the original data, resulting in consistency between each constraint-augmented instruction and the reformulated response.
 Specifically, we focus on the characteristics of responses that can be defined by rules, e.g., by checking the 220 distinct linguistic features in the LFTK package Lee and Lee (2023).
@@
 These diverse characteristics include punctuation-, word-, sentence-, paragraph-level occurrences, and frequencies.
-We construct rules constraining or specifying the characteristics of the original response.
+Thus we construct rules constraining or specifying the characteristics of the original response.
 In conjunction with the rules containing these characteristics, we also create rules specifying the format of responses to improve LLM’s format-following capability.
-To ensure that the rules selected or sampled for each sample are applicable and legal, we apply the following additional protocols: (1) The rules need to be applicable to the original sample.
+To ensure that the rules selected or sampled for each sample are applicable, we apply the following additional protocols: (1) The rules need to be applicable to the original sample.
 For instance, the rule “Generating a title before giving the response” is not applicable as we can not generate a title without the help of humans or other additional models.
-(2) The rules should not remove the content of the original response.
-For instance, the rule “Ensure the word xxx is not shown in the response” is not applicable since we can not directly remove this word from the response as it is part of the original response.
+(2) The rules should not include removing the content of the original response.
+For instance, the rule “Ensure the word xxx is not shown in the response” is not applicable since we can not directly remove this word from the response as the removal might disrupt the original semantic integrity.
 (3) The rules should be compatible with the original sample.
@@
 2.2.2 Rule Implementation
-To implement each rule as a program editing instructions and responses, we formulate each pre-defined rule as a tuple, $(\mathbf{S}_{k},f_{k},g_{k})$, where $\mathbf{S}_{k}$ represents the set of manually curated templates for creating instructions of the $k^{th}$ rule, while $f_{k}$ and $g_{k}$ are the corresponding functions to reformulate instructions and responses (if necessary), respectively.
+To implement each rule to original instructions and responses, we notate each pre-defined rule as a tuple for simplicity, $(\mathbf{S}_{k},f_{k},g_{k})$, where $\mathbf{S}_{k}$ represents the set of manually curated instruction templates for creating instructions of the $k^{th}$ rule, while $f_{k}$ and $g_{k}$ are the corresponding functions to reformulate instructions and responses (if necessary), respectively.
 Specifically, the function $f_{k}$ selects an appropriate template of a rule for a given data sample and augments the original instruction with an instruction generated by the template.
 It first selects a subset of rules applicable to the characteristics of the response (e.g., presence of keywords, number of sentences, etc.).
-Then, we randomly draw one rule out of the subset and create a formatted instruction of the rule from the template.
+Then, it randomly draw one rule out of the subset and create a formatted instruction of the rule from the template.
 Such rule instruction is appended to the original instruction.
@@
 For rules defining the case or format of certain parts in the response, modifications are needed.
-The detailed descriptions of each predefined rule can be found in Appendix C.
+The detailed descriptions of each predefined rule can be found in Appendix E.
 Specifically, the augmented response $y_{i,aug}$ will be optionally modified based on the selected rule $k$:
@@
 2.2.3 Multi-rule Implementation
-§ 2.2.2 illustrates how RuleR applies one rule to edit each data sample.
-To create more complex, diverse, and challenging samples, we can extend it to multiple rules randomly drawn from the feasible set of rules.
-We provide examples of multi-rule augmentation in Figure 3.
-It shows that the complexity of the augmented data grows dramatically, forcing the LLMs to learn to follow multiple constraints.
+To create more complex, diverse, and challenging samples, we can extend the previous process to multiple rules randomly drawn from the feasible set of rules.
+We provide examples of multi-rule augmentation in Appendix B, which forces LLMs to learn to follow multiple constraints.
+The detailed experimental setup can be found in Appendix A, including Implementation Details, Training Datasets, and Evaluation Metrics.
 Model Dataset Method Instruction Following Eval Pair-wise Open LLM Leaderboard
@@
 A, H, M, and T denote ARC, HellaSwag, MMLU, and TruthfulQA.
-3 Experimental Setup
-3.1 Implementation Details
+3 Experimental Results
+3.1 Main Results
+The main experimental results are presented in Table 1, containing the performance comparison on the Instruction Following Eval Zhou et al.
+(2023b), Pair-wise Comparison Winning Score, and the Open LLM Leaderboard Gao et al.
+(2021a), on 3 different base models and several different instruction tuning datasets.
+The Pair-wise Comparison Winning Score is calculated as (Num(Win)--Num(Lose))///Num(All) $+1$ and the values that are greater than $1.0$ represent better responses generated.
+Detailed descriptions of evaluation metrics can be found in the Appendix D.
+Compared to the baseline, our method has consistent improvements on the IF Eval benchmark, across different base models and datasets, which aims at measuring LLMs’ constraint-following abilities by using verifiable instructions.
+It is astonishing that our method can improve the IF Eval scores by approximately $10\%$ on some of the configurations, by just utilizing the rule-based recycling method on the original data, without any human/model edition.
+Moreover, the performances keep being positive on originally diverse and high-quality datasets like Recycled WizardLM Li et al.
+(2023a) and DEITA Liu et al.
+(2023), which further verify the potential of our method.
+Compared with existing methods which enhance LLMs’ constraint controllability by generating totally new data, our method focuses more on fully utilizing the potential of existing data.
+Furthermore, our method not only improves the constraint controllability but also keeps the general instruction-following ability of the original data.
+The Pair-wise comparison and Open LLM leaderboard results showcase comparable or sometimes better performances compared with the baseline models.
+We hypothesize that the additional constraints largely complicate the original instructions, as shown in Figure 3, thus forcing the LLMs to understand each constraint before generating responses, thus leading to potentially improved instruction-following abilities.
+3.2 Ablation Studies
+Evaluation Metrics IF Eval Pair-wise Open LLM
+Baseline 39.07 1.000 61.24
+Single Temp. 42.62 0.987 61.43
+Diverse Temp. (*) 47.46 1.044 61.05
+Max Rule = 1 46.14 1.168 61.22
+Max Rule = 2 47.09 1.117 61.15
+Max Rule = 3 (*) 47.46 1.044 61.05
+Max Rule = 4 46.89 1.003 61.55
+Max Rule = 5 44.36 1.013 60.06
+Aug Rate = 0.1 41.72 1.020 61.34
+Aug Rate = 0.3 42.08 1.037 61.20
+Aug Rate = 0.5 46.55 1.111 61.31
+Aug Rate = 0.7 46.23 1.111 61.18
+Aug Rate = 0.9 (*) 47.46 1.044 61.05
+Table 2: Ablation Study. “(*)” represents default.
+In this section, ablation experiments are conducted on Mistral-7B with the Alpaca-GPT4 dataset, aiming to evaluate the impact of several factors.
+Effect of Templates: As shown in Table 2, “Single Temp.” represents utilizing only one rule-instruction template for each rule, while “Diverse Temp.” represents utilizing several different templates, approximately $10$, with the same meaning for each rule and those templates are randomly sampled during the augmentation.
+“Diverse Temp.” demonstrates superior performance in both IF Eval and Pair-wise winning scores, with slightly lower accuracy on the Open LLM leaderboard.
+This result suggests that “Diverse Temp.” enhances the model’s constraint controllability while enhancing its general instruction-following capabilities compared to “Single Temp.” On the contrary, when fixing the rule templates to one single template, potential overfitting to the template might occur and thus negatively influence the model performances.
+Effect of Rule Numbers: “Max Rule = $x$” represents the setting in which at most $x$ different rules can be sampled and utilized on each original data sample.
+In the augmentation process, a random value will be sampled in the range of $[0,x]$ as the number of rules in this sample.
+The examples in Figure 2 showcase the scenario when only one rule is implemented and examples in Figure 3 showcase the scenario when multiple rules are implemented.
+Compared to the baseline, nearly all settings show performance improvements across the three evaluation metrics.
+However, the IF Eval score initially increases, reaching its peak when “Max Rule = $3$”, before declining.
+The Pair-wise score, on the other hand, consistently decreases from “Max Rule = $1$” to “Max Rule = $5$”.
+These results suggest that applying too many rules to a single sample may impair the LLM’s capability, even though sampling and applying multiple rules can be beneficial when done in moderation, which might be because the original instruction becomes so complex.
+Effect of Augmentation Rate: “Aug Rate = $x$” represents there is a probability of $x$ to apply our augmentation to each sample.
+It is observed that as the augmentation rate increases, performance improves on the IF Eval and mostly improves on the Pair-wise evaluation.
+This phenomenon indicates that increasing the augmentation rate primarily enhances the LLM’s constraint controllability, and it also has a positive impact on its general instruction-following capability.
+However, the gaps on IF Eval are much larger than the other 2 metrics, indicating this rate will mostly influence the multi-constraint controllability of LLM.
+Detailed Sub-Category Analysis: The detailed sub-category analysis can be found in Appendix C.
+4 Conclusion
+In this work, we proposed Rule-based Data Recycling (RuleR), which modified the original instructions and responses from an existing dataset by rule-defined constraints.
+The “recycled” data aims to enhance the LLMs’ capability to generate outputs fulfilling the constraints specified in the input, thereby improving the controllability of LLMs.
+RuleR took the first step of exploring rule-based data recycling, which can serve as a plug-and-play and easy-to-use method that converts any existing SFT datasets to new datasets for better controllability.
+Limitations
+Our method focuses on improving LLM controllability by rule-based editing of existing data, thereby avoiding the extra cost of data generation by humans or expert models.
+Though it saves the cost of human/model editing, the rules inevitably limit the types of constraints that can be applied to modify the original data.
+In the presented RuleR, all the constraints and rules are based on verifiable shallow syntactic characteristics such as the occurrence and frequency of words or sentences while lacking constraints and controllability on the semantic feature or content.
+This implies a potential of RuleR to be further enhanced by modifying the semantic content with a smaller model, which retains a comparable efficiency of rule-based editing.
+Appendix A Experimental Setup
+A.1 Implementation Details
 We utilize the prompt and code base from Vicuna Chiang et al. (2023a) and flash attention Dao et al. (2022) for all our experiments.
@@
 For training on Llama2-7B and Llama2-13B Touvron et al.
-(2023b), the maximum learning rate is set to $2\times 10^{-5}$ with a warmup rate of $0.03$ for $3$ epochs.
+(2023), the maximum learning rate is set to $2\times 10^{-5}$ with a warmup rate of $0.03$ for $3$ epochs.
 For training on Mistral-7B Jiang et al.
@@
 All other configurations are kept the same as the baselines.
-3.2 Training Datasets
-We utilize 5 SFT datasets to evaluate the effectiveness of our method: Alpaca dataset Taori et al.
-(2023), Alpaca-GPT4 dataset Peng et al.
-(2023), WizardLM dataset Xu et al.
-(2023), Recycled WizardLM Dataset Li et al.
-(2023a), DEITA dataset Liu et al.
-(2023).
-A detailed illustration of datasets can be found in the Appendix A.
-3.3 Evaluation Metrics
+A.2 Training Datasets
+We utilize 5 SFT datasets to evaluate the effectiveness of our method:
+Alpaca dataset Taori et al.
+(2023): This dataset consists of 52,000 instruction-following samples created using the self-instruct paradigm Wang et al.
+(2023b) and OpenAI’s text-davinci-003 model.
+Characterized as a classical dataset with moderate-quality attributes, it serves as the fundamental validation.
+Alpaca-GPT4 dataset Peng et al. (2023): This dataset is an enhanced Alpaca dataset that includes responses generated by GPT-4.
+WizardLM dataset Xu et al.
+(2023): This dataset is generated by the novel Evol-Instruct method, which utilizes ChatGPT-3.5 to rewrite instructions step by step into more complex ones and generate the corresponding responses.
+We utilize the 70k version in our method, which comprises 70,000 high-quality SFT samples.
+Recycled WizardLM Dataset Li et al.
+(2023a): This dataset is an improved version of the WizardLM dataset, by utilizing the Reflection-Tuning method.
+In the Reflection-Tuning, the initial dataset undergoes two main phases: Reflection on Instruction and Reflection on Response.
+In the first phase, specific criteria are carefully curated to evaluate and refine the initial instructions.
+During the second phase, responses are thoroughly examined and improved to align with the refined instructions.
+This process generates a dataset with superior quality compared to the original dataset.
+DEITA dataset Liu et al.
+(2023): This dataset leverages the DEITA (Data-Efficient Instruction Tuning for Alignment) method to select high-quality data from a pool comprised of several high-quality datasets, such as WizardLM and Alpaca.
+DEITA employs a score-first, diversity-aware data selection strategy to optimize the selection process.
+This strategy uses a GPT-as-a-judge scoring system that combines complexity and quality in a practical and straightforward manner.
+The scores are incorporated with the diversity-based selection, ensuring that all the data maintains high standards of complexity, quality, and diversity.
+A.3 Evaluation Metrics
 We employ 3 commonly accepted metrics for the evaluation, including IFEval (Instruction-Following Eval), Pair-wise Comparison, and Open LLM Leaderboard.
@@
 One model’s response can be noted as Win/Lose only if it has a Higher/Lower score in both directions.
-The specific evaluation prompt is presented in the Appendix B.
+The specific evaluation prompt is presented in the Appendix D.
 Open LLM Leaderboard employs Eval Harness Gao et al.
@@
 (2021b).
-4 Experimental Results
-4.1 Main Results
-The main experimental results are presented in Table 1, containing the performance comparison on the Instruction Following Eval Zhou et al.
-(2023b), Pair-wise Comparison Winning Score, and the Open LLM Leaderboard Gao et al.
-(2021a), on 3 different base models and several different instruction tuning datasets.
-The Pair-wise Comparison Winning Score is calculated as (Num(Win)--Num(Lose))///Num(All) $+1$ and the values that are greater than $1.0$ represent better responses generated.
-Detailed descriptions of evaluation metrics can be found in the Appendix B.
-Compared to the baseline, our method has consistent improvements on the IF Eval benchmark, across different base models and datasets, which aims at measuring LLMs’ constraint-following abilities by using verifiable instructions.
-It is astonishing that our method can improve the IF Eval scores by approximately $10\%$ on some of the configurations, by just utilizing the rule-based recycling method on the original data, without any human/model edition.
-Moreover, the performances keep being positive on originally diverse and high-quality datasets like Recycled WizardLM Li et al.
-(2023a) and DEITA Liu et al.
-(2023), which further verify the potential of our method.
-Compared with existing methods which enhance LLMs’ constraint controllability by generating totally new data, our method focuses more on fully utilizing the potential of existing data.
-Furthermore, our method not only improves the constraint controllability but also keeps the general instruction-following ability of the original data.
-The Pair-wise comparison and Open LLM leaderboard results showcase comparable or sometimes better performances compared with the baseline models.
-We hypothesize that the additional constraints largely complicate the original instructions, as shown in Figure 3, thus forcing the LLMs to understand each constraint before generating responses, thus leading to potentially improved instruction-following abilities.
-4.2 Ablation Studies
-Evaluation Metrics IF Eval Pair-wise Open LLM
-Baseline 39.07 1.000 61.24
-Single Temp. 42.62 0.987 61.43
-Diverse Temp. (*) 47.46 1.044 61.05
-Max Rule = 1 46.14 1.168 61.22
-Max Rule = 2 47.09 1.117 61.15
-Max Rule = 3 (*) 47.46 1.044 61.05
-Max Rule = 4 46.89 1.003 61.55
-Max Rule = 5 44.36 1.013 60.06
-Aug Rate = 0.1 41.72 1.020 61.34
-Aug Rate = 0.3 42.08 1.037 61.20
-Aug Rate = 0.5 46.55 1.111 61.31
-Aug Rate = 0.7 46.23 1.111 61.18
-Aug Rate = 0.9 (*) 47.46 1.044 61.05
-Table 2: Ablation Study. Average scores on “IF Eval” and “Open LLM” of Mistral-7B finetuned with ours-augmented Alpaca-GPT4 data. “(*)” represents default configuration.
+Appendix B Examples
+Examples of multi-rule augmentation are shown in Figure 3.
+[Instruction]
+Ensure the 2nd paragraph is enclosed in "-| |-" in the response.
+Write a story of no more than 150 words about a knight who defends their kingdom from a dragon.
+Make sure the 2nd sentence is enclosed within "\$ \$", the 4th sentence is enclosed within "\$\$\$ \$", the 6th sentence is enclosed within "## #", the 8th sentence is enclosed within "*** *" in the response.
+[Response]
+In the kingdom of Arvania, there lived Sir Cedric, a valiant knight, sworn to defend the land against any threat.
+\$One day, the kingdom came under attack from a fire-breathing dragon, terrorizing the people and burning homes.\$ The king called on Sir Cedric, charging him to defeat the dragon and restore peace to the land.
+\$\$\$-|Wielding his sword and shield, Sir Cedric rode out to face the dragon, fearlessly engaging it in combat.\$ The battle was fierce, with the dragon’s flames clashing against the knight’s steel.
+##But Sir Cedric was determined and unrelenting, his bravery and skill unmatched.|- #
+Finally, with a mighty blow, Sir Cedric struck the dragon down, ending its reign of terror.
+***The people of Arvania rejoiced, celebrating the bravery of their hero.* And Sir Cedric returned in triumph, hailed as the dragon-slaying knight.
+[Instruction]
+Ensure the 1st sentence is in uppercase, the 2nd sentence is in uppercase, the 3rd sentence is in uppercase in the response.
+Keep the word "way" bounded by "@ @" in the response if it appears.
+Describe an example of a time you used to influence in a positive way
+Keep the 2nd sentence wrapped in "[[ ]]", the 4th sentence wrapped in "*** ***", and the 5th sentence wrapped in "# #" in the response.
+[Response]
+I RECENTLY HAD A TEAM PROJECT AT WORK WHERE I HAD TO INFLUENCE MY TEAM MEMBERS TO COMPLETE THE TASK IN THE MOST EFFICIENT @WAY@ POSSIBLE.
+[[I IDENTIFIED AREAS WHERE OUR TEAM COULD IMPROVE, AND I CONSTANTLY COMMUNICATED AND ENCOURAGED MY TEAMMATES TO STAY FOCUSED AND ON TRACK.]] I ALSO SOUGHT FEEDBACK ON HOW WE COULD IMPROVE THE PROCESS.
+***Through my efforts, I managed to motivate my team to complete the project ahead of schedule while also maintaining high-quality work.*** #This helped to boost morale and foster collaboration between our team members.# Everyone was very satisfied with the resulting outcome.
+Figure 3: Examples with multiple rules selected and implemented.
+The randomly generated rule-instructions are colored in violet.
+The upper example is augmented by $2$ different rules (Paragraph Wrapping, and Sentence Wrapping); the bottom example is augmented by $3$ different rules (Sentence Case, Keyword Wrapping, and Sentence Wrapping).
+The format differences in new responses are highlighted in red.
+Appendix C Detailed Sub-Category Analysis
+In this section, detailed comparisons between our models and baseline models are provided in Table 3 for further analysis of the effects of our method.
+It is worth noting that the specific prompts used in IFEval are kept unknown in the construction of our rule templates, avoiding potential overfitting to the templates.
+Models trained with our recycled data outperform the baseline model consistently in “Case”, “Combination”, “Punctuation”, and “Start End” categories, which are partially well-recycled by our method.
+In the “Length” category, our models are only slightly better than the baseline model although our recycling method contains this kind of constraints.
+After further investigation, we find the performance in this category is mainly influenced by the original data length distributions.
+Since our method does not introduce more data samples, thus not able to improve the performance dramatically, but the better performance indeed provides the model with a better understanding of response length.
+Interestingly, the performance in the “Language” category also shows consistent improvement although we do not introduce any more data.
+Considering the consistency between this performance and the Pair-wise comparison performance, we hypothesize this improvement is caused by the better general instruction-following abilities provided by our method.
+The performance in the “Content” category presents one of the limitations of our Rule-based Recycling method, without utilizing other models or human experts to rewrite the instruction and response, it’s hard for our method to modify the content of the existing response.
+The performances of our models in the “Json” and “Keywords” categories are merely slightly lower, which is mainly affected by the diversity of original training datasets.
+Comparing the performance changes across augmentation rates, LLMs obtain better performances when the augmentation rate is higher, except for “Case”, indicating the easiness of case-related constraints for LLMs to understand and learn.
 Sub-Category Case Combination Content Json Keywords Language Length Punctuation Start End
@@
 The top section represents the performance by the strict criterion while the bottom represents the loose.
-In this section, ablation experiments are conducted on Mistral-7B with the Alpaca-GPT4 dataset.
-The experiments aim to evaluate the impact of several factors on our method: template diversity, the maximum number of rules utilized in each data sample, and the probability of applying augmentation to each sample.
-The performance of each configuration is evaluated on three evaluation metrics, Instruction Following Eval Pair-wise winning scores, and Open LLM Leaderboard, to simultaneously inspect the constraint controllability and general instruction-following ability.
-4.2.1 Effect of Templates
-This section illustrates the effectiveness of utilizing diverse rule templates for the augmentation.
-As shown in Table 2, “Single Temp.” represents utilizing only one rule-instruction template for each rule, while “Diverse Temp.” represents utilizing several different templates, approximately $10$, with the same meaning for each rule and those templates are randomly sampled during the augmentation.
-As indicated in the table, “Diverse Temp.” demonstrates superior performance in both IF Eval and Pair-wise winning scores, with slightly lower accuracy on the Open LLM leaderboard.
-This result suggests that “Diverse Temp.” enhances the model’s constraint controllability while enhancing its general instruction-following capabilities compared to “Single Temp.” On the contrary, when fixing the rule templates to one single template, potential overfitting to the template might occur and thus negatively influence the model performances.
-4.2.2 Effect of Rule Numbers
-This section illustrates the effects of implementing multiple random rules together in one single sample.
-“Max Rule = $x$” represents the setting in which at most $x$ different rules can be sampled and utilized on each original data sample.
-In the augmentation process, a random value will be sampled in the range of $[0,x]$ as the number of rules in this sample.
-The examples in Figure 2 showcase the scenario when only one rule is implemented and examples in Figure 3 showcase the scenario when multiple rules are implemented.
-Compared to the baseline, nearly all settings show performance improvements across the three evaluation metrics.
-However, as shown in the table, the IF Eval score initially increases, reaching its peak when “Max Rule = $3$”, before declining.
-The Pair-wise score, on the other hand, consistently decreases from “Max Rule = $1$” to “Max Rule = $5$”.
-Although no clear trend is observed on the Open LLM leaderboard, Max Rule = 5 shows the worst performance.
-These results suggest that applying too many rules to a single sample may impair the LLM’s capability, even though sampling and applying multiple rules can be beneficial when done in moderation, which might be because the original instruction becomes so complex for LLMs to learn.
-4.2.3 Effect of Augmentation Rate
-This section illustrates the effect of the augmentation rate in our method. “Aug Rate = $x$” represents there is a probability of $x$ to apply our augmentation to each sample.
-It is observed that as the augmentation rate increases, performance improves on the IF Eval and mostly improves on the Pair-wise evaluation.
-However, the results on the Open LLM leaderboard show limited improvement from the baseline, without a significant upward trend.
-This phenomenon indicates that increasing the augmentation rate primarily enhances the LLM’s constraint controllability, and it also has a positive impact on its general instruction-following capability.
-However, the gaps on IF Eval are much larger than the other 2 metrics, indicating this rate will mostly influence the multi-constraint controllability of LLM.
-4.3 Detailed Sub-Category Analysis
-In this section, detailed comparisons between our models and baseline models are provided in Table 3 for further analysis of the effects of our method.
-It is worth noting that the specific prompts used in IFEval are kept unknown in the construction of our rule templates, avoiding potential overfitting to the templates.
-Models trained with our recycled data outperform the baseline model consistently in “Case”, “Combination”, “Punctuation”, and “Start End” categories, which are partially well-recycled by our method.
-In the “Length” category, our models are only slightly better than the baseline model although our recycling method contains this kind of constraints.
-After further investigation, we find the performance in this category is mainly influenced by the original data length distributions.
-Since our method does not introduce more data samples, thus not able to improve the performance dramatically, but the better performance indeed provides the model with a better understanding of response length.
-Interestingly, the performance in the “Language” category also shows consistent improvement although we do not introduce any more data.
-Considering the consistency between this performance and the Pair-wise comparison performance, we hypothesize this improvement is caused by the better general instruction-following abilities provided by our method.
-The performance in the “Content” category presents one of the limitations of our Rule-based Recycling method, without utilizing other models or human experts to rewrite the instruction and response, it’s hard for our method to modify the content of the existing response.
-The performances of our models in the “Json” and “Keywords” categories are merely slightly lower, which is mainly affected by the diversity of original training datasets.
-Comparing the performance changes across augmentation rates, LLMs obtain better performances when the augmentation rate is higher, except for “Case”, indicating the easiness of case-related constraints for LLMs to understand and learn.
-5 Conclusion
-In this work, we proposed Rule-based Data Recycling (RuleR), which modified the original instructions and responses from an existing dataset by rule-defined constraints.
-The “recycled” data aims to enhance the LLMs’ capability to generate outputs fulfilling the constraints specified in the input, thereby improving the controllability of LLMs.
-RuleR took the first step of exploring rule-based data recycling.
-As a data generation approach, it is much cheaper than existing methods generating new data from scratch since RuleR does not rely on any human supervision or additional models.
-Our RuleR can serve as a plug-and-play and easy-to-use method that converts any existing instruction-tuning datasets to new datasets for gaining better controllability.
-Hence, our method is more accessible and widely applicable.
-RuleR provides a promising and sustainable strategy to continuously improve LLMs.
-Moreover, our method has great potential to be extended to other settings.
-For example, with the help of additional models, we can apply content-level or semantic-level rules to produce more diverse training data, as discussed in the next section.
-Limitations
-Our method focuses on improving LLM controllability by rule-based editing of existing data, thereby avoiding the extra cost of data generation by humans or expert models.
-Though it saves the cost of human/model editing, the rules inevitably limit the types of constraints that can be applied to modify the original data.
-In the presented RuleR, all the constraints and rules are based on verifiable shallow syntactic characteristics such as the occurrence and frequency of words or sentences while lacking constraints and controllability on the semantic feature or content.
-This implies a potential of RuleR to be further enhanced by modifying the semantic content with a smaller model, which retains a comparable efficiency of rule-based editing.
-Appendix A Training Dataset
-We utilize 5 SFT datasets to evaluate the effectiveness of our method:
-Alpaca dataset Taori et al.
-(2023): This dataset consists of 52,000 instruction-following samples created using the self-instruct paradigm Wang et al.
-(2023b) and OpenAI’s text-davinci-003 model.
-Characterized as a classical dataset with moderate-quality attributes, it serves as the fundamental validation.
-Alpaca-GPT4 dataset Peng et al. (2023): This dataset is an enhanced Alpaca dataset that includes responses generated by GPT-4.
-WizardLM dataset Xu et al.
-(2023): This dataset is generated by the novel Evol-Instruct method, which utilizes ChatGPT-3.5 to rewrite instructions step by step into more complex ones and generate the corresponding responses.
-We utilize the 70k version in our method, which comprises 70,000 high-quality SFT samples.
-Recycled WizardLM Dataset Li et al.
-(2023a): This dataset is an improved version of the WizardLM dataset, by utilizing the Reflection-Tuning method.
-In the Reflection-Tuning, the initial dataset undergoes two main phases: Reflection on Instruction and Reflection on Response.
-In the first phase, specific criteria are carefully curated to evaluate and refine the initial instructions.
-During the second phase, responses are thoroughly examined and improved to align with the refined instructions.
-This process generates a dataset with superior quality compared to the original dataset.
-DEITA dataset Liu et al.
-(2023): This dataset leverages the DEITA (Data-Efficient Instruction Tuning for Alignment) method to select high-quality data from a pool comprised of several high-quality datasets, such as WizardLM and Alpaca.
-DEITA employs a score-first, diversity-aware data selection strategy to optimize the selection process.
-This strategy uses a GPT-as-a-judge scoring system that combines complexity and quality in a practical and straightforward manner.
-The scores are incorporated with the diversity-based selection, ensuring that all the data maintains high standards of complexity, quality, and diversity.
-Appendix B Evaluation Metrics
+Appendix D Evaluation Metrics
 The prompt for pair-wise comparison is shown in Figure 4.
@@
 Figure 4: The prompt we used to request GPT4 to evaluate the responses.
-Appendix C Predefined Rules
+Appendix E Predefined Rules
 Rule Type Rule Name Example of an Instruction Template for the Rule